### v1

# The Enigma of Li-Rich Giants and its Relation with Temporal Variations Observed in Radial Velocity and Stellar Activity Signals

Abstract
Context. Despite the large number of studies focused on the characterization of Li-rich stars and understanding of the mechanisms leading to such enrichment, their origin remains a mystery.
Aims. Magnetic activity, in particular the phenomena usually associated with it, such as spots and plages, and the Li abundance (A(Li)) of stars, are in general thought to be connected. However, as of today it is unclear just how. In this work, we study a sample of young but evolved intermediate mass red giants, inhabitants of open clusters where planets have been searched. We aim at using radial velocity (RV) and stellar activity indicator signals to look for relations between Li abundances and stellar activity/variability.
Methods. We explore how the standard deviation (STD), peak to peak amplitude (PTP), mean and median of typical stellar activity indicators (BIS, FWHM, , and H index) change as a function of the Li content of 82 red giants. Furthermore, we compute weighted Pearson Correlation Coefficients () between time series of RV measurements and the stellar activity indicators for the stars in our sample. To aid our results, we also study Generalized Lomb-Scargle Periodograms (GLSP) to capture possible significant periodic temporal variations in our data.
Results. Our analysis indicates that the STD and PTP of BIS and FWHM, the mean and median of the H index and increase exponentially with A(Li) in our sample of red giants. Significant temporal variations and correlations between RVs and activity indicators also tend to be found preferentially for stars where high A(Li) is observed. Most of the Li-rich stars in our sample either show strong correlations of RV with at least one of the stellar activity indicators or reveal significant periodic temporal variations in their stellar activity indicators GLSPs consistent with those found for RV.
Key Words.:
red-giants – open clusters – lithium rich – stellar activity1 Introduction
Low () to intermediate mass stars () (Iben 1967a, b), before entering the red giant branch (RGB) of their evolution, go through a mixing episode called the first dredge up (FDU), which destroys most, if not all, of their surface lithium (Li). Stars as massive as the Sun, according to standard stellar evolution models, should not exhibit a surface lithium abundance, A(Li)111, exceeding 1.5 dex after the FDU. In general, observations concur this but 1% of observed giants (e.g Brown et al. 1989; Kumar et al. 2011; Liu et al. 2014; Casey et al. 2016; Kirby et al. 2016; Deepak & Reddy 2019; Gao et al. 2019; Pabst et al. 2020; Martell et al. 2021) present abnormally high A(Li) with the highest abundance observed to date in a red giant star being 5.62 dex (Kowkabany et al. 2022). These stars are known as Li-rich giants.
[…]
It is also generally accepted that higher than normal A(Li) is common among magnetically active stars, even if the reason for this is still poorly understood. This has been evidenced for different types of stars namely dwarfs (Favata et al. 1996; Zboril et al. 1997), magnetically peculiar stars (see Lyubimkov 2016 for a review), active binaries and pre-main-sequence stars (Pallavicini et al. 1992), active main sequence and zero-age main sequence stars (Xing et al. 2021) and active giant stars (Fekel & Balachandran 1994; Sneden et al. 2022).
In this work, we are interested in studying the possible connection between Li enhancement in red giant stars and their stellar activity. For this we start by looking for relations between several descriptive statistics of stellar activity indicators and A(Li) in a sample of red giant stars, some of which are lithium rich, where planets have been searched. These stars are inhabitants of open clusters, which is particularly favourable, as we are able to better constrain their evolutionary status. To strengthen our analysis we also look for correlations as well as periodic temporal variations in stellar activity signals that may connect variability observed in radial velocity (RV) to stellar activity. The present article is organized as follows, in Sect. 2 we provide an overview of our data. In Sect. 3 we explain the methods incurred in the data analysis. In Sect. 4 we discuss our results, and the main conclusions are laid out in Sect. 5.
2 Observations and Sample
Our sample is comprised of young but evolved intermediate mass red giants resident in 10 different open clusters (IC4651, IC4756, NGC2345, NGC2423, NGC3114, NGC3532, NGC3680, NGC4349, NGC5822 and NGC6705). These stars have been observed over the years with the HARPS spectograph (Mayor et al. 2003) at the ESO 3.6-metre telescope (R=115 000) and the CORALIE spectrograph (Queloz et al. 2000; Udry et al. 2000) at the Swiss 1.2-metre Leonhard Euler Telescope at ESO’s La Silla Observatory (R=60 000), as part of a program to search for exoplanets. For a complete description of the RV surveys see Lovis & Mayor (2007); Delgado Mena et al. (2018); Delgado Mena et al. (2023). The stellar parameters and Li abundances for this sample are presented in Tsantaki et al. (2023), obtained using the same spectra employed here.
In this work, the data available from CORALIE will not be treated. CORALIE data carries a much higher associated uncertainty in the measurements than HARPS data, due to its significantly lower resolution (R). Furthermore, CORALIE data was only collected for a subset of the stars in our sample, and the measurements for each star are few in number. For these reasons it is preferable to use only HARPS data.
Only stars with 5 or more observational data points were considered for analysis. This is such to increase the credibility of our results, while including the highest amount of observed stars possible. In total, 82 stars were analysed out of which 15 are canonically Li-rich, that is, show A(Li)1.5 dex. If a Population I star forming with a metallicity alike that of the Sun has in the beginning of its life A(Li)3.3 dex, that is, the approximate A(Li) of the interstellar medium, accounting for the dilution occurring during the FDU, in the red giant phase of its evolution it should be left with a maximum of A(Li)1.5 dex (e.g. Charbonnel & Balachandran 2000). Studying stars in open clusters, however, allows for a more precise estimate of this limit. Open clusters are ideal laboratories for studying Li enhancement given that they provide natural constraints on mass, composition, age and evolutionary state of their members. With such information it is possible to estimate the amount of Li that all stars in a given cluster should have at their current age using stellar evolution models. This allows us to better constrain what makes a star in a given cluster “Li-rich” (see Sun et al. 2022). Tsantaki et al. (2023) conducted such a study, providing estimations for the maximum A(Li) that stars in 32 different open clusters should have in their current evolutionary stage. Stars will be considered Li-rich according to the limits defined in Table 1 and the canonical definition. Special attention will also be given to stars that show high Li content in comparison with inhabitants of the same cluster in the same evolutionary stage, these will be deemed ‘Li-Rich-in-Context’ in accordance with Tsantaki et al. (2023) and Sun et al. (2022). All relevant stars come presented in Table 2. Some of these stars have been previously studied in the literature. A thorough discussion concerning the possible nature of the Li enrichment observed in these stars has been provided in Tsantaki et al. (2023), and is summarized in Table 2.
| Cluster | A(Li[dex] | [Fe/H] [dex] |
| IC4651 | 1.46 | -0.01 |
| IC4756 | 1.26 | -0.12 |
| NGC2345 | 1.37 | -0.22 |
| NGC2423 | 1.46 | -0.03 |
| NGC3114 | 1.46 | -0.10 |
| NGC3532 | 1.40 | -0.08 |
| NGC3680 | 1.26 | -0.15 |
| NGC4349 | 1.26 | -0.11 |
| NGC5822 | 1.46 | -0.10 |
| NGC6705 | 1.54(1) | 0.07 |
[…]
| IC4651No6333 | 1.21 | In Context | luminosity bump | 1.78 | 2.13 | - | Extra-Mixing |
| IC4651No9791 | 1.28 | In Context | luminosity bump | 1.75 | 1.85 | - | Extra-Mixing |
| IC4756No52 (∙) | 1.45 | A(Li) | RGB tip | 2.27 | 2.87 | - | Unknown |
| NGC2345No50 | 0.98 | In Context | RGB/early AGB | 5.84 | 5.27 | (¿)26(ξ) | Unknown/HBB |
| NGC2423No3 | 1.38 | In Context | RGB-Lum. bump? | 2.04 | 2.19 | - | Unknown |
| NGC3114No170(♯) | 1.50 | Canonical | red clump | 3.98 | 11.32 | 24 (♯♯) | Planet Accretion/Tidal Interactions |
| NGC3532No19 | 1.44 | A(Li) | red clump | 3.22 | 5.05 | 12(†) | Extra-Mixing(δ) |
| NGC3532No649(♭) | 3.27 | Canonical | - | 2.17 | - | 10(‡) | Extra-Mixing |
| NGC3532No670 | 1.47 | A(Li) | RGB tip | 3.05 | 4.64 | 20(†) | Unknown |
| NGC3680No13 | 1.25 | In Context | luminosity bump | 1.66 | 0.28 | 16 (ζ) | Extra-Mixing |
| NGC3680No53 | 1.21 | In Context | luminosity bump | 1.65 | 0.12 | 10 (ζ) | Extra-Mixing |
| NGC4349No9 | 1.26 | A(Li) | red clump | 3.00 | 9.58 | 16(∗) | Planet Accretion/Tidal Interactions |
| NGC4349No127 | 1.37 | A(Li) | RGB tip | 3.01 | 4.81 | 18(∗) | Unknown |
| NGC5822No102 | 1.34 | In Context | RGB base/red clump | 2.02 | 5.83 | - | Unknown |
| NGC5822No240 | 1.40 | In Context | RGB tip | 2.07 | 2.83 | 17† | Extra-Mixing |
| NGC6705No1101 | 1.53 | Canonical | RGB base | 3.67 | 11.24 | - | Li-rich Formation Environment |
| NGC6705No1117 | 1.57 | Canonical | red clump | 3.52 | 9.44 | - | Li-rich Formation Environment |
| NGC6705No1248 | 1.52 | Canonical | red clump | 3.46 | 8.43 | - | Li-rich Formation Environment |
| NGC6705No1256 | 1.59 | Canonical | RGB | 3.15 | 4.47 | - | Li-rich Formation Environment |
| NGC6705No1364 | 1.53 | Canonical | red clump | 3.67 | 12.24 | - | Li-rich Formation Environment |
| NGC6705No1658 | 1.51 | Canonical | red clump | 3.91 | 7.29 | - | Li-rich Formation Environment |
| NGC6705No411 | 1.57 | Canonical | RGB | 3.19 | 6.47 | - | Li-rich Formation Environment |
| NGC6705No660 | 1.53 | Canonical | red clump | 3.58 | 5.74 | - | Li-rich Formation Environment |
| NGC6705No669 | 1.51 | Canonical | red clump | 3.45 | 8.93 | - | Li-rich Formation Environment |
| NGC6705No779 | 1.57 | Canonical | RGB | 3.40 | 5.57 | - | Li-rich Formation Environment |
| NGC6705No816 | 1.53 | Canonical | red clump | 3.96 | 8.77 | - | Li-rich Formation Environment |
| NGC6705No916 | 1.53 | Canonical | red clump | 3.75 | 9.70 | - | Li-rich Formation Environment |
| NGC6705No963 | 1.54 | Canonical | red clump | 3.76 | 6.71 | - | Li-rich Formation Environment |
Notes: () This work; () Possible non-member: Katime Santrich et al. (2013). () As derived in Katime Santrich et al. (2013). () As derived in Smiljanic, R. et al. (2009). () As derived in Luck (1994). () Non-member: Mermilliod et al. (2008). () Possible non-member: Herzog et al. (1975); Frinchaboy & Majewski (2008); Bagdonas et al. (2018). () As derived in Holanda et al. (2022). () As derived in Holanda et al. (2019). () As derived in Peña Suárez et al. (2018).
3 Methods
To probe if the high A(Li) that we observe in red giant stars is connected to their stellar activity we explore two methods. First, we study how Li abundances relate to these five stellar activity indicators:
1.
the bisector inverse slope (BIS) of the cross correlation function;
2.
the full width at half maximum (FWHM) of the cross correlation function;
3.
the Hα index with both a 0.6 Å and 1.6 Å central bandpass (H06, H16);
4.
and the effective temperature (T).
In particular, we look for relations between A(Li) and the standard deviation (STD) and peak to peak amplitude (PTP) of all stellar activity indicators as well as the mean and median of H. Only statistically significant models at a 99% confidence level (with p-values of the F-test ) will be discussed.
Secondly, we would like to investigate if a possible connection between A(Li) and stellar activity is manifested in RV variability. So, we seek correlations between stellar activity indicators and RV measurements (see Figure 5). These will be computed according to the Pearson Correlation Coefficient, , which is here weighted by the uncertainty in the measurements. P-values for each correlation will also be provided to attest the statistical significance of the correlations. A p-value bellow 0.05 suggests that the correlation is significantly different from 0 at a 95% confidence level. In addition, periodic temporal variations of the RV measurements of these stars and the stellar activity indicators will be analysed by drawing Generalized Lomb-Scargle Periodograms (GLSP) (Zechmeister & Kürster 2009), to evaluate if they connect to the presence of Li.
When searching for planets it is important to devise methods to distinguish planetary signals from temporal variability that may arise from the phenomena appearing at the surface of the star as a consequence of stellar magnetic activity. The BIS and FWHM values of each observation are provided by the HARPS data reduction pipeline as part of the process to calculate the RV of a given spectrum. Correlations found between RV and respective BIS and FWHM measurements indicate that periodical variability noticed in RV signals may originate from stellar activity or stellar surface inhomogeneities, rather than planets. Variability observed in BIS and FWHM over time with approximately the same period as that observed in RV, can also further corroborate this. These types of studies have often been conducted to probe whether or not a planet really exists in orbit of a certain star (e.g. Queloz et al. 2001; Delgado Mena et al. 2018).
In studying how stellar activity impacts an RV signal, it is common to analyse spectral lines that serve as an activity proxy due to their well known correlations with the existence of active regions in the surface of the star. One such spectral line is the Hα line. Here, we determine Hα index values from spectra using ACTIN 2222https://github.com/gomesdasilva/ACTIN2 (Gomes da Silva et al. 2018, 2021). This tool to calculate stellar activity indices provides two measurements for the Hα index for two different central bandpasses: 0.6 Å and 1.6 Å (see Figure 6). Since it has been mostly used for dwarf stars (Gomes da Silva et al. 2022), there is a lack of an extensive study for stars as the ones in our sample, and it is not clear which central bandpass is optimal for the study of these stars. Recently, in analysing the strong RV signals of some of the stars in this sample (Delgado Mena et al. 2023), we found that H16 seems to be a more reliable activity indicator for these evolved stars. But this conclusion was based on a small sample. As such, and since the two bandpasses probe the stellar atmosphere at different layers being sensitive to different phenomena, in our analysis both are used.
We further extend our search for correlations and periodical variations to the Teff. Manifestations of magnetic activity, much like spots and plages are usually accompanied by temperature variations. Spots are cooler than their surrounding photosphere, while plages are hotter than their surrounding chromosphere. During periods of intense magnetic activity it is possible that, if the Teff is affected, the variations observed could correlate with the activity induced variations (e.g. Yana Galarza et al. 2019). We use the stellar spectral analysis package, FASMA333https://github.com/MariaTsantaki/FASMA-synthesis (Tsantaki et al. 2018), to determine the Teff from individual stellar spectra. FASMA relies on the radiative transfer code MOOG444https://www.as.utexas.edu/~chris/moog.html (Sneden 1973) to model the atmosphere and create a synthetic spectrum for a set of input parameters (surface gravity, ; iron metallicity, [Fe/H]; macro-turbulence velocity ; and projected rotational velocity ) of a given star. The synthetic spectrum is then statistically compared to the observed spectrum of the star and the parameters are iteratively adjusted using a non-linear least-squares algorithm to find the set that best fits observations. To capture reliable variations in Teff, when computing it, we fix all other input stellar parameters, and leave Teff to be computed freely for each individual stellar spectra. Stellar parameters for the stars in our sample have been derived in Tsantaki et al. (2023) also using FASMA.
4 Results and Discussion
As presented in Table 2 and in accordance with Mermilliod et al. (2008), the star NGC3532No649 is not a member of cluster NGC3532 making the calculation of its mass and evolutionary stage uncertain. For this reason, it will not be included in the following discussion despite it being Li-rich.
4.1 Is the Lithium Abundance of Red Giants related to Stellar Activity?
We start by analysing the dispersion (STD and PTP) of BIS, FWHM, T and H to probe potential connections between activity variability and A(Li). We also test if the mean and median of Hα, indicative of a star’s activity level, is related to A(Li). For the computation of these relations we only use stars with A(Li) as this is the limit of FASMA’s minimization algorithm in the determination of A(Li) (Tsantaki et al. 2023). Using these limit values would be a source of bias. We find that the STD and PTP of BIS and FWHM increase exponentially with A(Li) and the mean and median values of H06 also increase exponentially with A(Li). In particular, the vast majority of stars with high H06 index are Li enriched. On the other hand, we do not find A(Li) to be related to H16. These results suggest that the activity levels of the stars in our sample seem to increase with increasing A(Li). These results are in agreement with Sneden et al. (2022) who found that Li-rich giants are more chromospherically active when in comparison to their Li poor counterparts, as they are more likely to show strong He I10830 absorptions. We also find that correlates well with dispersion of BIS and FHWM as well as the mean and median of the H06 index. This is indicative that more magnetically active stars also tend to rotate faster. Gonçalves et al. (2020) also arrived at a similar conclusion. The best fit relations are listed in Table 3. Plots of the relations found for STD of BIS and FWHM as well as the mean of H06 are shown in Figure 1. The plots for PTP of BIS and FWHM and the median of H06 look similar to the ones shown in Figure 1 and are presented in Appendix B. Also in Appendix B are plots of the mean and median H16 as a function of A(Li), for which no clear relation was found. In addition to the results of the F-test, in Table 3 we also provide R2 values, with better fits having R2 closer to 1. R2 values for the relations we find are in general low as there is significant scatter around the lines of best fit.
| Formula | F-Statistic | p(F-Statistic) | R2 |
| STD(BIS)= | 20.62 | 2.2 | 0.22 |
| STD(FWHM)= | 13.59 | 0.0004 | 0.16 |
| PTP(BIS)= | 18.66 | 4.93 | 0.21 |
| PTP(FWHM)= | 11.73 | 0.001 | 0.14 |
| MeanHα06= | 11.55 | 0.001 | 0.14 |
| MedianHα06= | 12.12 | 0.0009 | 0.14 |
| = | 34.82 | 1.09 | 0.33 |
4.2 Manifestation of Stellar Activity in RV signals and its Connection to Lithium Enhancement
To further investigate these findings, we also looked at how RV variability correlates with the stellar activity indicators and whether this is linked to A(Li). All the correlations discussed here are shown in Figure 2 and listed in Appendix C. From our analysis, 22/81 stars show a strong correlation (0.4555Some of the analysed stars in our sample do not have a large number of data points. Because of this we chose to include moderate correlations which have the potential to become strong with the addition of new data. For ease of nomenclature we refer to all correlations 0.4 as strong.) between the RV variability and the BIS variability. The majority of the correlations were found to be negative. We noticed that 13/22 of the strong correlations found between RV and BIS appear for stars that are moderate and fast rotators, that is, show (e.g. de Medeiros et al. 1996; Tayar et al. 2015; Patton et al. 2023). This is likely a consequence of this diagnostic of line asymmetry losing sensitivity for slower rotators (e.g. Santos et al. 2003). There is, however, evidence in the literature (e.g. Carlberg et al. 2012; Sneden et al. 2022; Sayeed et al. 2023; Tsantaki et al. 2023) that there is a connection between Li enrichment in giants and . This is also suggested in our sample as we find that increases exponentially with A(Li) (shown in Appendix B and Table 3). Out of the 27 Li-rich stars listed in Table 2, 8 stars show strong correlations between RV and BIS.
In what concerns the FWHM, 20/81 stars show strong correlations with RV measurements. The correlations are preferentially positive. 6 of the Li-rich stars in our sample show strong correlations between RV and FWHM.
For the T we find that 15/81 stars show strong correlations with RV measurements. Correlations between the RV and Teff are mostly positive. Only 3 of the Li-rich stars show strong correlations between RV and T. There is a possibility that this is such because the individual spectra used to measure Teff values to look for variations may not have a signal to noise ratio high enough for this purpose.
As for H16, 20/81 stars show strong correlations with RV, which are mostly negative. 9 of the Li-rich stars show strong correlations between RV and H16.
Last but not least for H06, 19/81 stars show significant correlations with RV measurements, which are preferentially positive. We find that 10 of the Li-rich stars in our sample show strong correlations between RV and H06.
Overall, correlations were found for stars in various evolutionary stages and of different mass ranges as well. This is evidenced in the Hertzsprung–Russell (HR) diagram Figure 3, where different symbols indicate different numbers of strong correlations found per star, and the data points are sized according to stellar mass. In general, all symbols display a range of sizes and are distributed across the HR diagram revealing there is no preferred mass or age for correlations to be found. A summary of the significant correlations discussed can be found in Tables 4 and 5. Something to note from Table 5 is that inhabitants of cluster NGC6705, in general, show higher amounts of Li than the rest of the stars in our sample. These stars are particular as they also demonstrate unusually high amounts of -elements for their young age (e.g. Casamiquela et al. 2018). Messina et al. (2010) found that the G type stars of this cluster retain higher levels of activity up to the age of the cluster. We find that, 19/28 of these stars show at least one significant correlation with one of the parameters at study and 22/28 are fast rotators (see Table 5). So in general, the stars in this cluster are magnetically active and also rotate fast which could be connected to the abnormally high levels of Li observed.
In Figure 4 we show a violin plot of A(Li) as a function of the number of strong correlations observed per star. We may see that the density of points increases towards higher values of A(Li) with increasing number of correlations found. This is evidenced by the increasing width of the violins from left to right in high A(Li) regions and by the central tendency being slightly higher for stars that show 1 or more than 2 correlations between RV and stellar activity indicators. As such, strong correlations tend to appear more often for stars with higher amounts of Li in their atmospheres. Such a result also indicates that A(Li) seems to be connected with higher levels of stellar activity in our sample.
To complement the correlation analysis, GLSP of RV and stellar activity indicators were also plotted for all stars. GLSP are often used when analysing whether variability observed in RV signals is caused by the existence of a companion or by stellar activity. This is because, they allow us to determine if there is periodic variability occurring at the stellar surface that is equivalent to periodic variability observed in RV. It has been shown that magnetic phenomena appearing at the stellar surface can introduce variability in RV that is at a different phase relatively to variability observed in stellar activity indicator signals (e.g. Queloz et al. 2001; Santos et al. 2014; Delgado Mena et al. 2018). As such, when looking for correlations between RV and activity proxies we may obtain a weak correlation coefficient, but this does not necessarily mean that RV variability does not originate from stellar activity. Thus, finding peaks at similar periods in RV and again in different stellar activity signals with GLSP, is also an indicator that RV variability connects with stellar activity.
The results of the GLSP analysis are also presented in Tables 4 and 5. In summary, stars NGC2345No50, NGC2423No3, NGC3532No670, NGC4349No127 and NGC6705No1101 show significant peaks (above a false alarm probability (FAP) level of 1%) in their RV GLSPs which consistently appear around the same period for at least one of their stellar activity indicators. All relevant GLSPs are in Appendix D and Delgado Mena et al. (2023). All of these stars can be considered Li-rich (see Table 2) and most of them also present long-term and large amplitude RV variations which mimic the presence of planets but are of probable stellar origin (Delgado Mena et al. 2023). These stars with periodic RV variations lie on the RGB or RGB tip (they are among the most luminous in their cluster) rather than in the red clump, as found in recent works for super-Li-rich stars (Singh et al. 2019; Kumar et al. 2020). Interestingly, Jorissen et al. (2020) studying the binarity fraction of Li-rich giants found that some of the most luminous Li-rich giants show RV variability not caused by stellar companions, though no information about stellar activity is provided by the authors. In a similar way, Gonçalves et al. (2020) finds hints of RV variability in several Li-rich stars with detected longitudinal magnetic field, but the quantity of data does not allow to discern the real origin of such variability.
5 Conclusions
In summary based on the results from HARPS data, we find that the variability of BIS and FWHM and the mean and median of H06 increase exponentially with A(Li). This leads us to conclude that the stars in our sample that are more magnetically active also show higher amounts of Li in their atmosphere. We also find that stars with higher amounts of Li tend to show more correlations between RV and stellar activity indicators. The majority of stars, 19/27 (10/14 leaving aside the special cluster NGC6705), that can be considered Li-rich in this work, either show correlations of RV with at least one of the stellar activity indicators or show periodic variability in at least one of the stellar activity indicator signals with a similar period in RV. From the GLSP analysis we may also conclude that almost all the stars with strong RV variations that mimic the presence of planets are Li-rich (see also, Delgado Mena et al. 2016). This is particularly important when searching for exoplanets orbiting evolved stars, since our results show that detecting periodic RV variability in red giants with high A(Li) might be a signature of stellar activity instead of planets. It is interesting to note that Adamów et al. (2018) reported a higher than expected fraction of planetary companions around the Li-rich stars in their sample, pointing to a connection between Li enrichment and planetary presence/engulfment. However, the orbits they find are quite eccentric, something not expected if the RV were of stellar origin, and that supports the hypothesis of inner planets engulfment. The Li-rich stars with strong RV variability discussed here lie on the RGB or RGB tip, however, stars showing correlations between Li abundance and activity level are found all across the HR diagram. This might indicate that mild Li enrichment (since we do not have super-Li-rich giants in our sample) connected to stellar activity can occur at different evolutionary stages, but a large enhancement of Li is more particular and can only occur at certain evolutionary stages, such as, the He-flash (Singh et al. 2019; Kumar et al. 2020). It is still not clear then, whether the strong RV variability of mild Li-rich giants on the RGB is connected to their evolutionary stage and the possible super-Li rich phase to happen soon after (at the beginning of the He-core burning phase), or if it is a consequence of stellar activity.
Acknowledgements.
I.R. and E.D.M. acknowledge the support from Fundação para a Ciência e a Tecnologia (FCT) through national funds and from FEDER through COMPETE2020 by the following grants: UIDB/04434/2020 & UIDP/04434/2020. I.R. acknowledges the support from FCT through the grant 2022.03993.PTDC. E.D.M. acknowledges the support from FCT through Stimulus FCT contract 2021.01294.CEECIND. E.D.M and J.G.S. acknowledge the support from FCT through the grant 2022.04416.PTDC.References
- Adamów et al. (2018) Adamów, M., Niedzielski, A., Kowalik, K., et al. 2018, A&A, 613, A47
- Adamów et al. (2012) Adamów, M., Niedzielski, A., Villaver, E., Nowak, G., & Wolszczan, A. 2012, ApJ, 754, L15
- Adamów et al. (2014) Adamów, M., Niedzielski, A., Villaver, E., Wolszczan, A., & Nowak, G. 2014, A&A, 569, A55
- Aguilera-Gómez et al. (2016) Aguilera-Gómez, C., Chanamé, J., Pinsonneault, M. H., & Carlberg, J. K. 2016, ApJ, 829, 127
- Aguilera-Gómez et al. (2023) Aguilera-Gómez, C., Jones, M. I., & Chanamé, J. 2023, A&A, 670, A73
- Bagdonas et al. (2018) Bagdonas, V., Drazdauskas, A., Tautvaišienė, G., Smiljanic, R., & Chorniy, Y. 2018, Astronomy and Astrophysics, 615, A165
- Brown et al. (1989) Brown, J. A., Sneden, C., Lambert, D. L., & Dutchover, Edward, J. 1989, ApJS, 71, 293
- Cameron & Fowler (1971) Cameron, A. G. W. & Fowler, W. A. 1971, ApJ, 164, 111
- Cantiello & Langer (2010) Cantiello, M. & Langer, N. 2010, A&A, 521, A9
- Carlberg et al. (2012) Carlberg, J. K., Cunha, K., Smith, V. V., & Majewski, S. R. 2012, ApJ, 757, 109
- Casamiquela et al. (2018) Casamiquela, L., Carrera, R., Balaguer-Núñez, L., et al. 2018, A&A, 610, A66
- Casey et al. (2016) Casey, A. R., Ruchti, G., Masseron, T., et al. 2016, MNRAS, 461, 3336
- Charbonnel (1994) Charbonnel, C. 1994, A&A, 282, 811
- Charbonnel (1995) Charbonnel, C. 1995, ApJ, 453, L41
- Charbonnel & Balachandran (2000) Charbonnel, C. & Balachandran, S. C. 2000, arXiv preprint astro-ph/0005280
- de Medeiros et al. (1996) de Medeiros, J. R., Melo, C. H. F., & Mayor, M. 1996, A&A, 309, 465
- Deepak & Reddy (2019) Deepak & Reddy, B. E. 2019, MNRAS, 484, 2000
- Delgado Mena et al. (2023) Delgado Mena, E., Gomes da Silva, J., Faria, J. P., et al. 2023, A&A, 679, A94
- Delgado Mena et al. (2018) Delgado Mena, E., Lovis, C., Santos, N. C., et al. 2018, A&A, 619, A2
- Delgado Mena et al. (2016) Delgado Mena, E., Tsantaki, M., Sousa, S., et al. 2016, Astronomy & Astrophysics, 587, A66
- Denissenkov & Herwig (2004) Denissenkov, P. A. & Herwig, F. 2004, ApJ, 612, 1081
- Eggleton et al. (2008) Eggleton, P. P., Dearborn, D. S. P., & Lattanzio, J. C. 2008, ApJ, 677, 581
- Favata et al. (1996) Favata, F., Micela, G., & Sciortino, S. 1996, A&A, 311, 951
- Fekel & Balachandran (1994) Fekel, F. C. & Balachandran, S. 1994, in Cool Stars, Stellar Systems, and the Sun, Vol. 64, 279
- Frinchaboy & Majewski (2008) Frinchaboy, P. M. & Majewski, S. R. 2008, The Astronomical Journal, 136, 118
- Gao et al. (2019) Gao, Q., Shi, J.-R., Yan, H.-L., et al. 2019, ApJS, 245, 33
- Gomes da Silva et al. (2022) Gomes da Silva, J., Bensabat, A., Monteiro, T., & Santos, N. C. 2022, A&A, 668, A174
- Gomes da Silva et al. (2018) Gomes da Silva, J., Figueira, P., Santos, N., & Faria, J. 2018, Journal of Open Source Software, 3, 667
- Gomes da Silva et al. (2021) Gomes da Silva, J., Santos, N. C., Adibekyan, V., et al. 2021, A&A, 646, A77
- Gonçalves et al. (2020) Gonçalves, B. F. O., da Costa, J. S., de Almeida, L., Castro, M., & do Nascimento, J. D., J. 2020, MNRAS, 498, 2295
- Herzog et al. (1975) Herzog, A. D., Sanders, W. L., & Seggewiss, W. 1975, Astronomy and Astrophysics Supplement Series, 19, 211
- Holanda et al. (2019) Holanda, N., Pereira, C. B., & Drake, N. A. 2019, MNRAS, 482, 5275
- Holanda et al. (2022) Holanda, N., Ramos, A. A., Peña Suárez, V. J., Martinez, C. F., & Pereira, C. B. 2022, Monthly Notices of the Royal Astronomical Society, 516, 4484
- Iben (1967a) Iben, Icko, J. 1967a, ApJ, 147, 650
- Iben (1967b) Iben, Icko, J. 1967b, ApJ, 147, 624
- Jorissen et al. (2020) Jorissen, A., Van Winckel, H., Siess, L., et al. 2020, A&A, 639, A7
- Katime Santrich et al. (2013) Katime Santrich, O. J., Pereira, C. B., & Drake, N. A. 2013, Astronomy and Astrophysics, 554, A2
- Kirby et al. (2016) Kirby, E. N., Guhathakurta, P., Zhang, A. J., et al. 2016, 819, 135
- Kowkabany et al. (2022) Kowkabany, J., Ezzeddine, R., Charbonnel, C., et al. 2022, arXiv e-prints, arXiv:2209.02184
- Kumar et al. (2020) Kumar, Y. B., Reddy, B. E., Campbell, S. W., et al. 2020, Nature Astronomy, 4, 1059
- Kumar et al. (2011) Kumar, Y. B., Reddy, B. E., & Lambert, D. L. 2011, ApJ, 730, L12
- Lagarde et al. (2024) Lagarde, N., Minkevičiūtė, R., Drazdauskas, A., et al. 2024, A&A, 684, A70
- Liu et al. (2014) Liu, Y. J., Tan, K. F., Wang, L., et al. 2014, ApJ, 785, 94
- Lovis & Mayor (2007) Lovis, C. & Mayor, M. 2007, Astronomy and Astrophysics, 472, 657
- Luck (1994) Luck, R. E. 1994, The Astrophysical Journal Supplement Series, 91, 309
- Lyubimkov (2016) Lyubimkov, L. S. 2016, Astrophysics, 59, 411
- Martell et al. (2021) Martell, S. L., Simpson, J. D., Balasubramaniam, A. G., et al. 2021, MNRAS, 505, 5340
- Mayor et al. (2003) Mayor, M., Pepe, F., Queloz, D., et al. 2003, The Messenger, 114, 20
- Mermilliod et al. (2008) Mermilliod, J. C., Mayor, M., & Udry, S. 2008, Astronomy and Astrophysics, 485, 303
- Messina et al. (2010) Messina, S., Parihar, P., Koo, J.-R., et al. 2010, A&A, 513, A29
- Pabst et al. (2020) Pabst, C. H. M., Goicoechea, J. R., Teyssier, D., et al. 2020, A&A, 639, A2
- Palacios et al. (2004) Palacios, A., Charbonnel, C., Talon, S., & Siess, L. 2004, arXiv e-prints, astro
- Pallavicini et al. (1992) Pallavicini, R., Randich, S., & Giampapa, M. 1992, Astronomy and Astrophysics, 253, 185
- Patton et al. (2023) Patton, R. A., Pinsonneault, M. H., Cao, L., et al. 2023, arXiv e-prints, arXiv:2303.08151
- Peña Suárez et al. (2018) Peña Suárez, V. J., Sales Silva, J. V., Katime Santrich, O. J., Drake, N. A., & Pereira, C. B. 2018, ApJ, 854, 184
- Queloz et al. (2001) Queloz, D., Henry, G. W., Sivan, J. P., et al. 2001, A&A, 379, 279
- Queloz et al. (2000) Queloz, D., Mayor, M., Weber, L., et al. 2000, Astronomy and Astrophysics, 354, 99
- Randich et al. (2020) Randich, S., Pasquini, L., Franciosini, E., et al. 2020, A & A, 640, L1
- Romano et al. (2021) Romano, D., Magrini, L., Randich, S., et al. 2021, A & A, 653, A72
- Santos et al. (2014) Santos, N. C., Mortier, A., Faria, J. P., et al. 2014, A&A, 566, A35
- Santos et al. (2003) Santos, N. C., Udry, S., Mayor, M., et al. 2003, A&A, 406, 373
- Sayeed et al. (2023) Sayeed, M., Ness, M. K., Montet, B. T., et al. 2023, arXiv e-prints, arXiv:2306.03323
- Singh et al. (2019) Singh, R., Reddy, B. E., Bharat Kumar, Y., & Antia, H. M. 2019, ApJ, 878, L21
- Smiljanic, R. et al. (2009) Smiljanic, R., Gauderon, R., North, P., et al. 2009, A&A, 502, 267
- Sneden et al. (2022) Sneden, C., Afşar, M., Bozkurt, Z., et al. 2022, ApJ, 940, 12
- Sneden (1973) Sneden, C. A. 1973, PhD thesis, University of Texas
- Sun et al. (2022) Sun, Q., Deliyannis, C. P., Twarog, B. A., et al. 2022, Monthly Notices of the Royal Astronomical Society, 513, 5387
- Tayar et al. (2015) Tayar, J., Ceillier, T., García-Hernández, D. A., et al. 2015, ApJ, 807, 82
- Tsantaki et al. (2018) Tsantaki, M., Andreasen, D. T., Teixeira, G. D. C., et al. 2018, MNRAS, 473, 5066
- Tsantaki et al. (2023) Tsantaki, M., Delgado-Mena, E., Bossini, D., et al. 2023, A&A, 674, A157
- Udry et al. (2000) Udry, S., Mayor, M., Naef, D., et al. 2000, Astronomy and Astrophysics, 356, 590
- Xing et al. (2021) Xing, L.-F., Li, Y.-C., Chang, L., Wang, C.-J., & Bai, J.-M. 2021, A&A, 653, A28
- Yana Galarza et al. (2019) Yana Galarza, J., Meléndez, J., Lorenzo-Oliveira, D., et al. 2019, MNRAS, 490, L86
- Zboril et al. (1997) Zboril, M., Byrne, P. B., & Rolleston, W. R. J. R. 1997, Monthly Notices of the Royal Astronomical Society, 284, 685
- Zechmeister & Kürster (2009) Zechmeister, M. & Kürster, M. 2009, A&A, 496, 577
| Star | (RV,BIS) | (RV,FWHM) | (RV,Teff) | (RV,H1.6Å) | (RV,H0.6Å) | A(Li) [dex] | Mass (M⊙) | (km/s) | n |
| IC4651No10393 | -0.500 | 1.801 | 1.870 | 23 | |||||
| IC4651No11453 | -0.007 | 2.699 | 1.750 | 33 | |||||
| IC4651No12935 | 0.315 | 1.800 | 0.000 | 5 | |||||
| IC4651No14527 | -0.500 | 1.867 | 1.810 | 31 | |||||
| IC4651No6333 | 1.209 | 1.779 | 2.130 | 5 | |||||
| IC4651No7646 | -0.500 | 2.438 | 0.000 | 31 | |||||
| IC4651No8540 | 0.715 | 1.797 | 0.000 | 27 | |||||
| IC4651No9025 | 0.621 | 1.815 | 0.000 | 30 | |||||
| IC4651No9122 | -0.500 | 1.795 | 0.680 | 74 | |||||
| IC4651No9791 | 1.277 | 1.750 | 1.850 | 33 | |||||
| IC4756No14 | -0.016 | 1.973 | 2.030 | 5 | |||||
| IC4756No38 | 0.783 | 2.005 | 0.000 | 7 | |||||
| IC4756No44 | 0.823 | 2.016 | 2.830 | 6 | |||||
| IC4756No52 | 1.447 | 2.272 | 2.870 | 16 | |||||
| NGC2345No50 | 0.977 | 5.843 | 5.270 | 26 | |||||
| NGC2423No3 | 1.377 | 2.035 | 2.190 | 60 | |||||
| NGC2423No56 | 1.035 | 1.991 | 0.000 | 7 | |||||
| NGC3114No150 | 1.340 | 4.490 | 16.190 | 11 | |||||
| NGC3114No170 | 1.503 | 3.981 | 11.320 | 6 | |||||
| NGC3114No181 | 0.309 | 4.022 | 3.950 | 11 | |||||
| NGC3114No223 | -0.500 | 2.508 | 0.330 | 8 | |||||
| NGC3114No238 | 0.533 | 3.807 | 4.090 | 10 | |||||
| NGC3114No262 | 1.190 | 3.884 | 9.220 | 31 | |||||
| NGC3114No283 | 1.415 | 4.663 | 12.470 | 30 | |||||
| NGC3114No6 | 1.092 | 4.612 | 7.480 | 44 | |||||
| NGC3532No100 | 0.952 | 3.379 | 4.870 | 19 | |||||
| NGC3532No122 | 1.344 | 2.873 | 9.570 | 20 | |||||
| NGC3532No160 | 1.125 | 3.149 | 4.780 | 8 | |||||
| NGC3532No19 | 1.443 | 3.218 | 5.050 | 46 | |||||
| NGC3532No221 | 0.186 | 4.818 | 5.890 | 22 | |||||
| NGC3532No522 | 1.202 | 4.412 | 12.480 | 6 | |||||
| NGC3532No596 | 1.153 | 3.126 | 7.210 | 31 | |||||
| NGC3532No670 | 1.475 | 3.047 | 4.640 | 29 | |||||
| NGC3680No13 | 1.252 | 1.659 | 0.280 | 7 | |||||
| NGC3680No26 | 1.192 | 1.704 | 0.960 | 18 | |||||
| NGC3680No34 | 0.246 | 1.752 | 3.010 | 12 | |||||
| NGC3680No41 | 0.633 | 1.641 | 0.000 | 18 | |||||
| NGC3680No44 | -0.500 | 1.690 | 2.550 | 6 | |||||
| NGC3680No53 | 1.211 | 1.652 | 0.120 | 18 | |||||
| NGC4349No127 | 1.375 | 3.007 | 4.810 | 56 | |||||
| NGC4349No168 | 1.000 | 3.358 | 5.510 | 35 | |||||
| NGC4349No174 | 0.843 | 3.003 | 3.190 | 22 | |||||
| NGC4349No203 | 1.105 | 3.296 | 5.310 | 10 | |||||
| NGC4349No5 | -0.500 | 3.134 | 7.270 | 35 | |||||
| NGC4349No9 | 1.262 | 3.005 | 9.580 | 24 | |||||
| NGC5822No102 | 1.348 | 2.020 | 5.830 | 5 | |||||
| NGC5822No1 | 0.126 | 2.274 | 3.120 | 8 | |||||
| NGC5822No201 | 1.009 | 2.422 | 1.600 | 23 | |||||
| NGC5822No240 | 1.401 | 2.071 | 2.830 | 7 | |||||
| NGC5822No316 | 0.636 | 2.205 | 0.000 | 7 | |||||
| NGC5822No375 | -0.076 | 2.142 | 2.190 | 19 | |||||
| NGC5822No443 | -0.138 | 2.111 | 1.830 | 8 | |||||
| NGC5822No8 | 0.431 | 2.262 | 0.360 | 36 |
| Star | (RV,BIS) | (RV,FWHM) | (RV,Teff) | (RV,H1.6Å) | (RV,H0.6Å) | A(Li) [dex] | Mass (M⊙) | (km/s) | n |
| NGC6705No1090 | 1.380 | 3.800 | 6.050 | 9 | |||||
| NGC6705No1101 | 1.527 | 3.674 | 11.240 | 25 | |||||
| NGC6705No1111 | 1.425 | 3.618 | 7.110 | 10 | |||||
| NGC6705No1117 | 1.567 | 3.521 | 9.440 | 10 | |||||
| NGC6705No1145 | 1.385 | 3.355 | 7.300 | 11 | |||||
| NGC6705No1184 | 0.215 | 3.448 | 4.420 | 11 | |||||
| NGC6705No1248 | 1.525 | 3.455 | 8.430 | 9 | |||||
| NGC6705No1256 | 1.591 | 3.146 | 4.470 | 14 | |||||
| NGC6705No1286 | 1.400 | 3.709 | 9.940 | 21 | |||||
| NGC6705No1364 | 1.530 | 3.670 | 12.240 | 13 | |||||
| NGC6705No136 | 1.082 | 3.792 | 3.730 | 7 | |||||
| NGC6705No1423 | 1.455 | 3.852 | 5.480 | 13 | |||||
| NGC6705No1446 | 1.194 | 3.451 | 7.140 | 6 | |||||
| NGC6705No160 | 1.477 | 3.740 | 6.950 | 22 | |||||
| NGC6705No1625 | 1.088 | 3.087 | 4.100 | 19 | |||||
| NGC6705No1658 | 1.512 | 3.910 | 7.290 | 28 | |||||
| NGC6705No1837 | 1.495 | 3.493 | 10.520 | 9 | |||||
| NGC6705No2000 | 1.443 | 4.040 | 4.760 | 12 | |||||
| NGC6705No320 | 1.492 | 3.876 | 5.500 | 8 | |||||
| NGC6705No411 | 1.573 | 3.193 | 6.470 | 8 | |||||
| NGC6705No660 | 1.533 | 3.577 | 5.740 | 7 | |||||
| NGC6705No669 | 1.511 | 3.451 | 8.930 | 6 | |||||
| NGC6705No686 | 1.467 | 3.727 | 9.480 | 6 | |||||
| NGC6705No779 | 1.569 | 3.397 | 5.570 | 22 | |||||
| NGC6705No816 | 1.533 | 3.956 | 8.770 | 6 | |||||
| NGC6705No827 | -0.087 | 4.040 | 4.800 | 8 | |||||
| NGC6705No916 | 1.532 | 3.745 | 9.700 | 9 | |||||
| NGC6705No963 | 1.542 | 3.757 | 6.710 | 9 |
Appendix A Methods for Data Analysis
Appendix B Statistically Significant Relations
Appendix C Correlation Tables
| Star | (RV,BIS) | p(BIS) | (RV,FWHM) | p(FWHM) | (RV,Teff) | p(Teff) | (RV,H1.6Å) | p(H1.6Å) | (RV,H0.6Å) | p(H0.6Å) | A(Li) [dex] | Mass (M⊙) | (km/s) | n |
| IC4651No10393 | -0.12 | 0.58 | -0.03 | 0.89 | -0.11 | 0.63 | -0.25 | 0.26 | -0.32 | 0.14 | -0.500 | 1.801 | 1.870 | 23 |
| IC4651No11453 | -0.09 | 0.62 | 0.06 | 0.75 | 0.17 | 0.33 | 0.01 | 0.95 | -0.001 | 0.996 | -0.007 | 2.699 | 1.750 | 33 |
| IC4651No12935 | -0.02 | 0.98 | 0.25 | 0.68 | 0.11 | 0.86 | -0.44 | 0.45 | 0.07 | 0.91 | 0.315 | 1.800 | 0.000 | 5 |
| IC4651No14527 | -0.02 | 0.90 | 0.03 | 0.88 | 0.23 | 0.22 | -0.29 | 0.12 | -0.34 | 0.06 | -0.500 | 1.867 | 1.810 | 31 |
| IC4651No6333 | -0.81 | 0.10 | 0.28 | 0.64 | -0.13 | 0.84 | -0.65 | 0.23 | -0.23 | 0.71 | 1.209 | 1.779 | 2.130 | 5 |
| IC4651No7646 | -0.01 | 0.95 | 0.29 | 0.13 | 0.02 | 0.90 | -0.13 | 0.49 | -0.18 | 0.34 | -0.500 | 2.438 | 0.000 | 31 |
| IC4651No8540 | -0.28 | 0.15 | 0.24 | 0.23 | -0.01 | 0.97 | -0.24 | 0.24 | 0.23 | 0.25 | 0.715 | 1.797 | 0.000 | 27 |
| IC4651No9025 | -0.23 | 0.24 | 0.09 | 0.66 | -0.12 | 0.55 | 0.20 | 0.33 | -0.22 | 0.26 | 0.621 | 1.815 | 0.000 | 30 |
| IC4651No9122 | 0.18 | 0.13 | -0.12 | 0.33 | -0.05 | 0.66 | 0.10 | 0.41 | 0.19 | 0.11 | -0.500 | 1.795 | 0.680 | 74 |
| IC4651No9791 | -0.32 | 0.07 | 0.03 | 0.85 | -0.07 | 0.70 | 0.07 | 0.69 | 0.07 | 0.71 | 1.277 | 1.750 | 1.850 | 33 |
| IC4756No14 | -0.14 | 0.82 | 0.69 | 0.19 | -0.65 | 0.24 | 0.06 | 0.92 | 0.25 | 0.68 | -0.016 | 1.973 | 2.030 | 5 |
| IC4756No38 | -0.42 | 0.26 | -0.10 | 0.79 | 0.01 | 0.98 | -0.29 | 0.45 | 0.31 | 0.42 | 0.783 | 2.005 | 0.000 | 7 |
| IC4756No44 | 0.46 | 0.35 | 0.06 | 0.90 | -0.16 | 0.77 | -0.23 | 0.67 | 0.51 | 0.30 | 0.823 | 2.016 | 2.830 | 6 |
| IC4756No52 | 0.04 | 0.90 | -0.08 | 0.78 | -0.26 | 0.35 | 0.92 | 1.54E-06 | -0.08 | 0.77 | 1.447 | 2.272 | 2.870 | 16 |
| NGC2345No50 | 0.31 | 0.12 | -0.51 | 0.01 | 0.02 | 0.91 | 0.60 | 1.07E-03 | 0.56 | 2.86E-03 | 0.977 | 5.843 | 5.270 | 26 |
| NGC2423No3 | 0.56 | 3.40E-06 | 0.28 | 0.03 | 0.11 | 0.40 | 0.03 | 0.79 | -0.07 | 0.61 | 1.377 | 2.035 | 2.190 | 60 |
| NGC2423No56 | -0.11 | 0.82 | 0.12 | 0.79 | -0.07 | 0.88 | -0.91 | 4.38E-03 | 0.04 | 0.93 | 1.035 | 1.991 | 0.000 | 7 |
| NGC3114No150 | -0.52 | 0.10 | -0.06 | 0.86 | 0.19 | 0.58 | -0.17 | 0.61 | -0.24 | 0.47 | 1.340 | 4.490 | 16.190 | 11 |
| NGC3114No170 | -0.48 | 0.34 | -0.54 | 0.27 | 0.25 | 0.64 | -0.62 | 0.19 | -0.62 | 0.19 | 1.503 | 3.981 | 11.320 | 6 |
| NGC3114No181 | -0.23 | 0.47 | 0.62 | 0.03 | -0.59 | 0.04 | 0.40 | 0.20 | 0.26 | 0.42 | 0.309 | 4.022 | 3.950 | 11 |
| NGC3114No223 | -0.19 | 0.64 | -0.23 | 0.59 | -0.49 | 0.22 | 0.17 | 0.69 | 0.07 | 0.87 | -0.500 | 2.508 | 0.330 | 8 |
| NGC3114No238 | 0.26 | 0.44 | 0.26 | 0.44 | -0.01 | 0.97 | 0.29 | 0.38 | 0.15 | 0.65 | 0.533 | 3.807 | 4.090 | 10 |
| NGC3114No262 | -0.48 | 0.01 | -0.03 | 0.89 | 0.29 | 0.11 | -0.04 | 0.82 | 0.16 | 0.39 | 1.190 | 3.884 | 9.220 | 31 |
| NGC3114No283 | 0.49 | 0.01 | 0.20 | 0.30 | 0.09 | 0.64 | 0.06 | 0.74 | 0.27 | 0.15 | 1.415 | 4.663 | 12.470 | 30 |
| NGC3114No6 | -0.28 | 0.06 | -0.42 | 4.81E-03 | -0.18 | 0.23 | 0.35 | 0.02 | 0.32 | 0.03 | 1.092 | 4.612 | 7.480 | 44 |
| NGC3532No100 | 0.05 | 0.85 | 0.20 | 0.41 | -0.14 | 0.56 | 0.19 | 0.43 | 0.09 | 0.71 | 0.952 | 3.379 | 4.870 | 19 |
| NGC3532No122 | -0.60 | 4.94E-03 | 0.14 | 0.56 | 0.09 | 0.70 | 0.08 | 0.74 | 0.14 | 0.54 | 1.344 | 2.873 | 9.570 | 20 |
| NGC3532No160 | -0.75 | 0.03 | 0.53 | 0.18 | -0.65 | 0.08 | 0.56 | 0.15 | 0.19 | 0.65 | 1.125 | 3.149 | 4.780 | 8 |
| NGC3532No19 | -0.12 | 0.44 | -0.07 | 0.66 | 0.11 | 0.47 | 0.20 | 0.17 | -0.11 | 0.45 | 1.443 | 3.218 | 5.050 | 46 |
| NGC3532No221 | -0.41 | 0.05 | -0.27 | 0.20 | 0.62 | 1.58E-03 | -0.41 | 0.05 | 0.10 | 0.66 | 0.186 | 4.818 | 5.890 | 22 |
| NGC3532No522 | -0.61 | 0.20 | 0.79 | 0.06 | 0.43 | 0.39 | 0.02 | 0.97 | 0.10 | 0.85 | 1.202 | 4.412 | 12.480 | 6 |
| NGC3532No596 | 0.08 | 0.66 | 0.01 | 0.95 | 0.13 | 0.49 | -0.60 | 3.15E-04 | 0.10 | 0.61 | 1.153 | 3.126 | 7.210 | 31 |
| NGC3532No649 | -0.26 | 0.08 | 0.10 | 0.49 | 0.06 | 0.69 | 0.13 | 0.36 | 0.16 | 0.26 | 3.271 | 2.167 | 0.000 | 49 |
| NGC3532No670 | 0.18 | 0.34 | -0.13 | 0.49 | 0.09 | 0.63 | 0.35 | 0.06 | 0.42 | 0.02 | 1.475 | 3.047 | 4.640 | 29 |
| NGC3680No13 | 0.83 | 0.02 | 0.08 | 0.86 | 0.58 | 0.18 | 0.45 | 0.31 | 0.39 | 0.38 | 1.252 | 1.659 | 0.280 | 7 |
| NGC3680No26 | 0.02 | 0.94 | 0.06 | 0.82 | 0.17 | 0.50 | 0.28 | 0.27 | 0.35 | 0.15 | 1.192 | 1.704 | 0.960 | 18 |
| NGC3680No34 | -0.13 | 0.68 | -0.38 | 0.22 | 0.02 | 0.95 | -0.12 | 0.71 | -0.15 | 0.64 | 0.246 | 1.752 | 3.010 | 12 |
| NGC3680No41 | -0.03 | 0.91 | -0.15 | 0.55 | 0.20 | 0.42 | 0.16 | 0.53 | 0.01 | 0.97 | 0.633 | 1.641 | 0.000 | 18 |
| NGC3680No44 | -0.02 | 0.98 | 0.79 | 0.06 | 0.00 | 0.99 | 0.77 | 0.08 | 0.46 | 0.36 | -0.500 | 1.690 | 2.550 | 6 |
| NGC3680No53 | -0.06 | 0.81 | -0.04 | 0.88 | 0.31 | 0.20 | -0.03 | 0.91 | -0.10 | 0.70 | 1.211 | 1.652 | 0.120 | 18 |
| NGC4349No127 | 0.04 | 0.76 | -0.03 | 0.80 | -0.07 | 0.59 | 0.26 | 0.05 | 0.11 | 0.44 | 1.375 | 3.007 | 4.810 | 56 |
| NGC4349No168 | -0.17 | 0.34 | 0.13 | 0.46 | -0.04 | 0.81 | -0.07 | 0.69 | 0.03 | 0.85 | 1.000 | 3.358 | 5.510 | 35 |
| NGC4349No174 | -0.25 | 0.25 | -0.51 | 0.02 | 0.50 | 0.02 | 0.07 | 0.76 | 0.08 | 0.72 | 0.843 | 3.003 | 3.190 | 22 |
| NGC4349No203 | 0.24 | 0.50 | -0.47 | 0.17 | 0.15 | 0.69 | 0.27 | 0.45 | 0.21 | 0.56 | 1.105 | 3.296 | 5.310 | 10 |
| NGC4349No5 | -0.24 | 0.16 | -0.37 | 0.03 | 0.37 | 0.03 | -0.04 | 0.84 | -0.05 | 0.79 | -0.500 | 3.134 | 7.270 | 35 |
| NGC4349No9 | -0.48 | 0.02 | 0.37 | 0.07 | -0.04 | 0.86 | 0.13 | 0.53 | 0.57 | 3.24E-03 | 1.262 | 3.005 | 9.580 | 24 |
| NGC5822No102 | -0.84 | 0.07 | 1.00 | 1.20E-09 | 0.33 | 0.59 | -0.68 | 0.20 | -0.68 | 0.21 | 1.348 | 2.020 | 5.830 | 5 |
| NGC5822No1 | -0.20 | 0.63 | 0.10 | 0.81 | -0.05 | 0.91 | -0.02 | 0.96 | 0.03 | 0.95 | 0.126 | 2.274 | 3.120 | 8 |
| NGC5822No201 | -0.13 | 0.56 | -0.21 | 0.34 | -0.38 | 0.07 | -0.12 | 0.58 | -0.09 | 0.68 | 1.009 | 2.422 | 1.600 | 23 |
| NGC5822No240 | -0.11 | 0.82 | -0.26 | 0.58 | -0.16 | 0.73 | -0.20 | 0.67 | -0.20 | 0.67 | 1.401 | 2.071 | 2.830 | 7 |
| NGC5822No316 | -0.41 | 0.36 | 0.12 | 0.80 | -0.53 | 0.22 | 0.59 | 0.16 | 0.40 | 0.37 | 0.636 | 2.205 | 0.000 | 7 |
| NGC5822No375 | -0.21 | 0.37 | -0.07 | 0.76 | -0.35 | 0.13 | -0.39 | 0.09 | -0.19 | 0.41 | -0.076 | 2.142 | 2.190 | 19 |
| NGC5822No443 | -0.35 | 0.40 | 0.43 | 0.29 | -0.31 | 0.46 | 0.26 | 0.54 | -0.20 | 0.64 | -0.138 | 2.111 | 1.830 | 8 |
| NGC5822No8 | -0.061 | 0.724 | 0.151 | 0.379 | -0.100 | 0.563 | 0.119 | 0.490 | 0.042 | 0.807 | 0.431 | 2.262 | 0.360 | 36 |
| NGC6705No1090 | 0.01 | 0.98 | 0.23 | 0.56 | -0.08 | 0.84 | -0.08 | 0.84 | 0.11 | 0.78 | 1.380 | 3.800 | 6.050 | 9 |
| NGC6705No1101 | -0.40 | 0.05 | 0.22 | 0.29 | 0.12 | 0.57 | -0.15 | 0.46 | 0.13 | 0.53 | 1.527 | 3.674 | 11.240 | 25 |
| NGC6705No1111 | -0.28 | 0.43 | -0.48 | 0.16 | 0.27 | 0.45 | 0.55 | 0.10 | 0.02 | 0.95 | 1.425 | 3.618 | 7.110 | 10 |
| NGC6705No1117 | -0.25 | 0.49 | -0.13 | 0.72 | 0.22 | 0.54 | -0.57 | 0.09 | -0.47 | 0.17 | 1.567 | 3.521 | 9.440 | 10 |
| NGC6705No1145 | -0.56 | 0.07 | 0.68 | 0.02 | 0.39 | 0.23 | -0.02 | 0.96 | 0.28 | 0.40 | 1.385 | 3.355 | 7.300 | 11 |
| NGC6705No1184 | -0.07 | 0.84 | 0.14 | 0.69 | 0.43 | 0.19 | -0.50 | 0.11 | -0.25 | 0.45 | 0.215 | 3.448 | 4.420 | 11 |
| NGC6705No1248 | 0.09 | 0.82 | 0.44 | 0.24 | 0.27 | 0.48 | -0.50 | 0.17 | -0.29 | 0.45 | 1.525 | 3.455 | 8.430 | 9 |
| NGC6705No1256 | 0.43 | 0.12 | -0.15 | 0.61 | 0.28 | 0.34 | 0.11 | 0.71 | -0.32 | 0.27 | 1.591 | 3.146 | 4.470 | 14 |
| NGC6705No1286 | -0.71 | 3.99E-04 | -0.35 | 0.13 | 0.11 | 0.63 | 0.20 | 0.39 | 0.67 | 1.32E-03 | 1.400 | 3.709 | 9.940 | 21 |
| NGC6705No1364 | -0.11 | 0.73 | 0.31 | 0.30 | 0.05 | 0.87 | 0.26 | 0.38 | 0.32 | 0.28 | 1.530 | 3.670 | 12.240 | 13 |
| NGC6705No136 | 0.05 | 0.91 | 0.10 | 0.83 | 0.47 | 0.29 | -0.17 | 0.71 | -0.44 | 0.32 | 1.082 | 3.792 | 3.730 | 7 |
| NGC6705No1423 | -0.31 | 0.31 | 0.08 | 0.79 | 0.15 | 0.63 | 0.25 | 0.41 | -0.15 | 0.63 | 1.455 | 3.852 | 5.480 | 13 |
| NGC6705No1446 | -0.02 | 0.97 | 0.34 | 0.51 | 0.67 | 0.14 | -0.58 | 0.23 | 0.83 | 0.04 | 1.194 | 3.451 | 7.140 | 6 |
| NGC6705No160 | -0.29 | 0.19 | 0.29 | 0.19 | 0.22 | 0.33 | -0.11 | 0.62 | 0.11 | 0.63 | 1.477 | 3.740 | 6.950 | 22 |
| NGC6705No1625 | -0.44 | 0.06 | -0.11 | 0.66 | 0.25 | 0.31 | -0.20 | 0.40 | 0.41 | 0.08 | 1.088 | 3.087 | 4.100 | 19 |
| NGC6705No1658 | -0.26 | 0.19 | -0.06 | 0.76 | 0.07 | 0.72 | -0.10 | 0.63 | 0.13 | 0.50 | 1.512 | 3.910 | 7.290 | 28 |
| NGC6705No1837 | -0.77 | 0.01 | 0.52 | 0.15 | 0.61 | 0.08 | -0.03 | 0.94 | -0.06 | 0.88 | 1.495 | 3.493 | 10.520 | 9 |
| NGC6705No2000 | -0.13 | 0.70 | -0.05 | 0.88 | 0.30 | 0.35 | 0.29 | 0.36 | -0.44 | 0.16 | 1.443 | 4.040 | 4.760 | 12 |
| NGC6705No320 | 0.01 | 0.98 | -0.44 | 0.27 | -0.21 | 0.62 | 0.51 | 0.20 | 0.60 | 0.12 | 1.492 | 3.876 | 5.500 | 8 |
| NGC6705No411 | 0.07 | 0.86 | 0.61 | 0.11 | -0.10 | 0.82 | -0.47 | 0.24 | 0.52 | 0.18 | 1.573 | 3.193 | 6.470 | 8 |
| NGC6705No660 | -0.21 | 0.65 | -0.62 | 0.14 | 0.47 | 0.29 | -0.11 | 0.82 | -0.43 | 0.34 | 1.533 | 3.577 | 5.740 | 7 |
| NGC6705No669 | 0.38 | 0.46 | -0.25 | 0.64 | 0.13 | 0.80 | 0.14 | 0.79 | 0.48 | 0.34 | 1.511 | 3.451 | 8.930 | 6 |
| NGC6705No686 | -0.05 | 0.93 | 0.35 | 0.50 | 0.15 | 0.78 | -0.07 | 0.89 | -0.36 | 0.48 | 1.467 | 3.727 | 9.480 | 6 |
| NGC6705No779 | 0.34 | 0.13 | -0.32 | 0.15 | -0.18 | 0.43 | 0.18 | 0.43 | -0.05 | 0.83 | 1.569 | 3.397 | 5.570 | 22 |
| NGC6705No816 | -0.19 | 0.72 | 0.10 | 0.85 | 0.41 | 0.41 | -0.05 | 0.93 | -0.58 | 0.23 | 1.533 | 3.956 | 8.770 | 6 |
| NGC6705No827 | 0.06 | 0.89 | -0.54 | 0.22 | 0.23 | 0.62 | 0.24 | 0.61 | -0.03 | 0.95 | -0.087 | 4.040 | 4.800 | 8 |
| NGC6705No916 | -0.58 | 0.13 | -0.09 | 0.83 | 0.12 | 0.77 | -0.07 | 0.86 | 0.21 | 0.62 | 1.532 | 3.745 | 9.700 | 9 |
| NGC6705No963 | -0.28 | 0.47 | 0.27 | 0.48 | 0.03 | 0.93 | -0.21 | 0.60 | 0.14 | 0.71 | 1.542 | 3.757 | 6.710 | 9 |
Appendix D Generalized Lomb-Scargle Periodograms
The remainder of the GLPs mentioned in the text can be found in Delgado Mena et al. (2023).

### v2

@@
 Abstract
-Context. Despite the large number of studies focused on the characterization of Li-rich stars and understanding of the mechanisms leading to such enrichment, their origin remains a mystery.
+Context. Despite the large number of studies focused on the characterisation of Li-rich stars and understanding the mechanisms leading to such enrichment, their origin remains a mystery.
 Aims.
-Magnetic activity, in particular the phenomena usually associated with it, such as spots and plages, and the Li abundance (A(Li)) of stars, are in general thought to be connected.
-However, as of today it is unclear just how.
-In this work, we study a sample of young but evolved intermediate mass red giants, inhabitants of open clusters where planets have been searched.
-We aim at using radial velocity (RV) and stellar activity indicator signals to look for relations between Li abundances and stellar activity/variability.
+Magnetic activity, particularly the phenomena usually associated with it (e.g.
+spots and plages), and the Li abundance (A(Li)) of stars, are in general thought to be connected.
+As of today, however, just how they are connected is unclear.
+In this work, we study a sample of young but evolved intermediate-mass red giants that are inhabitants of open clusters where planets have been searched for.
+Our aim is to use radial velocity (RV) and stellar activity indicator signals to look for relations between Li abundances and stellar activity or variability.
 Methods.
-We explore how the standard deviation (STD), peak to peak amplitude (PTP), mean and median of typical stellar activity indicators (BIS, FWHM, , and H index) change as a function of the Li content of 82 red giants.
-Furthermore, we compute weighted Pearson Correlation Coefficients () between time series of RV measurements and the stellar activity indicators for the stars in our sample.
-To aid our results, we also study Generalized Lomb-Scargle Periodograms (GLSP) to capture possible significant periodic temporal variations in our data.
+We explored how the standard deviation (STD), peak-to-peak amplitude (PTP), mean, and median of typical stellar activity indicators (BIS, FWHM, , and H index) change as a function of the Li content of 82 red giants.
+Furthermore, we computed weighted Pearson correlation coefficients () between time series of RV measurements and the stellar activity indicators for the stars in our sample.
+To aid our results, we also studied generalized Lomb-Scargle periodograms (GLSP) to capture possible significant periodic temporal variations in our data.
 Results.
-Our analysis indicates that the STD and PTP of BIS and FWHM, the mean and median of the H index and increase exponentially with A(Li) in our sample of red giants.
+Our analysis indicates that the STD and PTP of BIS and FWHM, the mean and median of the H index, and increase exponentially with A(Li) in our sample of red giants.
 Significant temporal variations and correlations between RVs and activity indicators also tend to be found preferentially for stars where high A(Li) is observed.
-Most of the Li-rich stars in our sample either show strong correlations of RV with at least one of the stellar activity indicators or reveal significant periodic temporal variations in their stellar activity indicators GLSPs consistent with those found for RV.
+Most of the Li-rich stars in our sample either show strong correlations of RV with at least one of the stellar activity indicators or reveal significant periodic temporal variations in their GLSPs of stellar activity indicators that are consistent with those found for RV.
 Key Words.:
 red-giants – open clusters – lithium rich – stellar activity1 Introduction
-Low () to intermediate mass stars () (Iben 1967a, b), before entering the red giant branch (RGB) of their evolution, go through a mixing episode called the first dredge up (FDU), which destroys most, if not all, of their surface lithium (Li).
+Before entering the red giant branch (RGB) of their evolution, low- () to intermediate-mass stars () (Iben 1967a, b) go through a mixing episode called the first dredge up (FDU), which destroys most, if not all, of their surface lithium (Li).
 Stars as massive as the Sun, according to standard stellar evolution models, should not exhibit a surface lithium abundance, A(Li)111, exceeding 1.5 dex after the FDU.
-In general, observations concur this but 1% of observed giants (e.g Brown et al.
+In general, observations concur with this, but 1% of observed giants (e.g.
+Brown et al.
 1989; Kumar et al.
@@
 2020; Martell et al.
-2021) present abnormally high A(Li) with the highest abundance observed to date in a red giant star being 5.62 dex (Kowkabany et al.
+2021) present abnormally high A(Li), with the highest abundance observed to date in a red giant star being 5.62 dex (Kowkabany et al.
 2022).
@@
 It is also generally accepted that higher than normal A(Li) is common among magnetically active stars, even if the reason for this is still poorly understood.
-This has been evidenced for different types of stars namely dwarfs (Favata et al.
+This has been shown in different types of stars, including dwarfs (Favata et al.
 1996; Zboril et al.
-1997), magnetically peculiar stars (see Lyubimkov 2016 for a review), active binaries and pre-main-sequence stars (Pallavicini et al.
-1992), active main sequence and zero-age main sequence stars (Xing et al.
-2021) and active giant stars (Fekel & Balachandran 1994; Sneden et al.
+1997); magnetically peculiar stars (see Lyubimkov 2016 for a review); active binaries and pre-main-sequence stars (Pallavicini et al.
+1992); active main-sequence and zero-age main-sequence stars (Xing et al.
+2021); and active giant stars (Fekel & Balachandran 1994; Sneden et al.
 2022).
 In this work, we are interested in studying the possible connection between Li enhancement in red giant stars and their stellar activity.
-For this we start by looking for relations between several descriptive statistics of stellar activity indicators and A(Li) in a sample of red giant stars, some of which are lithium rich, where planets have been searched.
+As such, we started by looking for relations between several descriptive statistics of stellar activity indicators and A(Li) in a sample of red giant stars, some of which are lithium rich, where planets have been searched.
 These stars are inhabitants of open clusters, which is particularly favourable, as we are able to better constrain their evolutionary status.
-To strengthen our analysis we also look for correlations as well as periodic temporal variations in stellar activity signals that may connect variability observed in radial velocity (RV) to stellar activity.
-The present article is organized as follows, in Sect.
-2 we provide an overview of our data.
+To strengthen our analysis, we also looked for correlations as well as periodic temporal variations in stellar activity signals that may connect the variability observed in radial velocity (RV) to stellar activity.
+The present article is organised as follows: In Sect.
+2, we provide an overview of our data.
 In Sect.
-3 we explain the methods incurred in the data analysis.
+3, we explain the methods incurred in the data analysis.
 In Sect.
-4 we discuss our results, and the main conclusions are laid out in Sect.
+4, we discuss our results, and the main conclusions are laid out in Sect.
 5.
 2 Observations and sample
-Our sample is comprised of young but evolved intermediate mass red giants resident in 10 different open clusters (IC4651, IC4756, NGC2345, NGC2423, NGC3114, NGC3532, NGC3680, NGC4349, NGC5822 and NGC6705).
-These stars have been observed over the years with the HARPS spectograph (Mayor et al.
+Our sample is composed of young but evolved intermediate-mass red giants residing in ten different open clusters (IC4651, IC4756, NGC2345, NGC2423, NGC3114, NGC3532, NGC3680, NGC4349, NGC5822, and NGC6705).
+These stars have been observed over the years with the HARPS spectrograph (Mayor et al.
 2003) at the ESO 3.6-metre telescope (R=115 000) and the CORALIE spectrograph (Queloz et al.
 2000; Udry et al.
-2000) at the Swiss 1.2-metre Leonhard Euler Telescope at ESO’s La Silla Observatory (R=60 000), as part of a program to search for exoplanets.
-For a complete description of the RV surveys see Lovis & Mayor (2007); Delgado Mena et al.
+2000) at the Swiss 1.2-metre Leonhard Euler Telescope at ESO’s La Silla Observatory (R=60 000) as part of a programme to search for exoplanets.
+(For a complete description of the RV surveys see Lovis & Mayor (2007); Delgado Mena et al.
 (2018); Delgado Mena et al.
-(2023).
-The stellar parameters and Li abundances for this sample are presented in Tsantaki et al.
-(2023), obtained using the same spectra employed here.
-In this work, the data available from CORALIE will not be treated.
-CORALIE data carries a much higher associated uncertainty in the measurements than HARPS data, due to its significantly lower resolution (R).
-Furthermore, CORALIE data was only collected for a subset of the stars in our sample, and the measurements for each star are few in number.
-For these reasons it is preferable to use only HARPS data.
-Only stars with 5 or more observational data points were considered for analysis.
-This is such to increase the credibility of our results, while including the highest amount of observed stars possible.
-In total, 82 stars were analysed out of which 15 are canonically Li-rich, that is, show A(Li)1.5 dex.
-If a Population I star forming with a metallicity alike that of the Sun has in the beginning of its life A(Li)3.3 dex, that is, the approximate A(Li) of the interstellar medium, accounting for the dilution occurring during the FDU, in the red giant phase of its evolution it should be left with a maximum of A(Li)1.5 dex (e.g.
+(2023).) The stellar parameters and Li abundances for this sample are presented in Tsantaki et al.
+(2023), who used the same spectra employed here.
+In this work, the data available from CORALIE are not treated.
+CORALIE data carry a much higher associated uncertainty in the measurements than HARPS data due to a significantly lower resolution (R).
+Furthermore, CORALIE data were only collected for a subset of stars in our sample, and the measurements for each star are few in number.
+For these reasons we preferred to use only HARPS data.
+Only stars with five or more observational data points were considered for analysis.
+This limitation increases the credibility of our results while including the most observed stars as possible.
+In total, 82 stars were analysed out of which 15 are canonically Li-rich, that is they show A(Li)1.5 dex.
+If a Population I star forming with a metallicity similar to that of the Sun has in the beginning of its life A(Li)3.3 dex, that is, the approximate A(Li) of the interstellar medium, accounting for the dilution occurring during the FDU, it should be left with a maximum of A(Li)1.5 dex in the red giant phase of its evolution (e.g.
 Charbonnel & Balachandran 2000).
 Studying stars in open clusters, however, allows for a more precise estimate of this limit.
-Open clusters are ideal laboratories for studying Li enhancement given that they provide natural constraints on mass, composition, age and evolutionary state of their members.
-With such information it is possible to estimate the amount of Li that all stars in a given cluster should have at their current age using stellar evolution models.
-This allows us to better constrain what makes a star in a given cluster “Li-rich” (see Sun et al.
+Open clusters are ideal laboratories for studying Li enhancement given that they provide natural constraints on the mass, composition, age, and evolutionary state of their members.
+With such information, it is possible to estimate the amount of Li that all stars in a given cluster should have at their current age using stellar evolution models.
+This allows us to better constrain what makes a star in a given cluster ‘Li-rich’ (see Sun et al.
 2022).
@@
 (2023) conducted such a study, providing estimations for the maximum A(Li) that stars in 32 different open clusters should have in their current evolutionary stage.
-Stars will be considered Li-rich according to the limits defined in Table 1 and the canonical definition.
-Special attention will also be given to stars that show high Li content in comparison with inhabitants of the same cluster in the same evolutionary stage, these will be deemed ‘Li-Rich-in-Context’ in accordance with Tsantaki et al.
+Stars are considered Li-rich according to the limits defined in Table 2 and the canonical definition.
+Special attention is given to stars that show high Li content in comparison to inhabitants of the same cluster in the same evolutionary stage.
+We deem such stars ‘Li-Rich-in-Context’, in accordance with Tsantaki et al.
 (2023) and Sun et al.
 (2022).
-All relevant stars come presented in Table 2.
-Some of these stars have been previously studied in the literature.
+All relevant stars are presented in Table 3.
+Some of these stars have been studied previously in the literature.
 A thorough discussion concerning the possible nature of the Li enrichment observed in these stars has been provided in Tsantaki et al.
-(2023), and is summarized in Table 2.
+(2023), and it is summarised in Table 3.
 | Cluster | A(Li[dex] | [Fe/H] [dex] |
@@
 | NGC6705No963 | 1.54 | Canonical | red clump | 3.76 | 6.71 | - | Li-rich Formation Environment |
-Notes: () This work; () Possible non-member: Katime Santrich et al.
+() This work; () Possible non-member: Katime Santrich et al.
 (2013).
@@
 3 Methods
-To probe if the high A(Li) that we observe in red giant stars is connected to their stellar activity we explore two methods.
-First, we study how Li abundances relate to these five stellar activity indicators:
+To probe if the high A(Li) that we observe in red giant stars is connected to their stellar activity, we explored two methods.
+First, we studied how Li abundances relate to the following stellar activity indicators:
 1.
@@
 and the effective temperature (T).
-In particular, we look for relations between A(Li) and the standard deviation (STD) and peak to peak amplitude (PTP) of all stellar activity indicators as well as the mean and median of H.
-Only statistically significant models at a 99% confidence level (with p-values of the F-test ) will be discussed.
-Secondly, we would like to investigate if a possible connection between A(Li) and stellar activity is manifested in RV variability.
-So, we seek correlations between stellar activity indicators and RV measurements (see Figure 5).
-These will be computed according to the Pearson Correlation Coefficient, , which is here weighted by the uncertainty in the measurements.
-P-values for each correlation will also be provided to attest the statistical significance of the correlations.
-A p-value bellow 0.05 suggests that the correlation is significantly different from 0 at a 95% confidence level.
-In addition, periodic temporal variations of the RV measurements of these stars and the stellar activity indicators will be analysed by drawing Generalized Lomb-Scargle Periodograms (GLSP) (Zechmeister & Kürster 2009), to evaluate if they connect to the presence of Li.
-When searching for planets it is important to devise methods to distinguish planetary signals from temporal variability that may arise from the phenomena appearing at the surface of the star as a consequence of stellar magnetic activity.
+In particular, we looked for relations between A(Li) and the standard deviation (STD) and peak-to-peak amplitude (PTP) of all stellar activity indicators as well as the mean and median of H.
+Only statistically significant models at a 99% confidence level (with p-values of the F-test ) are discussed.
+Secondly, we investigated if a possible connection between A(Li) and stellar activity manifests in RV variability.
+So, we sought correlations between stellar activity indicators and RV measurements (see Figure 5).
+These were computed according to the Pearson correlation coefficient, , which is weighted here by the uncertainty in the measurements.
+The p-values for each correlation are also provided to attest to the statistical significance of the correlations.
+A p-value bellow 0.05 suggests that the correlation is significantly different from zero at a 95% confidence level.
+In addition, periodic temporal variations of the RV measurements of these stars and the stellar activity indicators were analysed by drawing generalised Lomb-Scargle Periodograms (GLSPs (Zechmeister & Kürster 2009)) in order to evaluate if they connect to the presence of Li.
+When searching for planets, it is important to devise methods to distinguish planetary signals from temporal variability that may arise from the phenomena appearing at the surface of a star as a consequence of stellar magnetic activity.
 The BIS and FWHM values of each observation are provided by the HARPS data reduction pipeline as part of the process to calculate the RV of a given spectrum.
-Correlations found between RV and respective BIS and FWHM measurements indicate that periodical variability noticed in RV signals may originate from stellar activity or stellar surface inhomogeneities, rather than planets.
-Variability observed in BIS and FWHM over time with approximately the same period as that observed in RV, can also further corroborate this.
-These types of studies have often been conducted to probe whether or not a planet really exists in orbit of a certain star (e.g.
+Correlations found between RV and respective BIS and FWHM measurements indicate that periodical variability noticed in RV signals may originate from stellar activity or stellar surface inhomogeneities rather than planets.
+Variability observed in BIS and FWHM over time with approximately the same period as that observed in RV can also further corroborate this.
+These types of studies have often been conducted to probe whether or not a planet really exists in the orbit of a certain star (e.g.
 Queloz et al.
@@
 2018).
-In studying how stellar activity impacts an RV signal, it is common to analyse spectral lines that serve as an activity proxy due to their well known correlations with the existence of active regions in the surface of the star.
+In studying how stellar activity impacts an RV signal, it is common to analyse spectral lines that serve as an activity proxy due to their well-known correlations with the existence of active regions on the surface of a star.
 One such spectral line is the Hα line.
-Here, we determine Hα index values from spectra using ACTIN 2222https://github.com/gomesdasilva/ACTIN2 (Gomes da Silva et al.
+Here, we determined Hα index values from spectra using ACTIN 2444https://github.com/gomesdasilva/ACTIN2 (Gomes da Silva et al.
 2018, 2021).
@@
 Since it has been mostly used for dwarf stars (Gomes da Silva et al.
-2022), there is a lack of an extensive study for stars as the ones in our sample, and it is not clear which central bandpass is optimal for the study of these stars.
+2022), there is a lack of an extensive study for stars such as the ones in our sample, and it is not clear which central bandpass is optimal for the study of these stars.
 Recently, in analysing the strong RV signals of some of the stars in this sample (Delgado Mena et al.
@@
 But this conclusion was based on a small sample.
-As such, and since the two bandpasses probe the stellar atmosphere at different layers being sensitive to different phenomena, in our analysis both are used.
-We further extend our search for correlations and periodical variations to the Teff.
+As such, and since the two bandpasses probe the stellar atmosphere at different layers that are sensitive to different phenomena, we used both bandpasses in our analysis.
+We further extended our search for correlations and periodical variations to the Teff.
 Manifestations of magnetic activity, much like spots and plages are usually accompanied by temperature variations.
 Spots are cooler than their surrounding photosphere, while plages are hotter than their surrounding chromosphere.
-During periods of intense magnetic activity it is possible that, if the Teff is affected, the variations observed could correlate with the activity induced variations (e.g.
+During periods of intense magnetic activity, it is possible that if the Teff is affected, the variations observed could correlate with the activity-induced variations (e.g.
 Yana Galarza et al.
 2019).
-We use the stellar spectral analysis package, FASMA333https://github.com/MariaTsantaki/FASMA-synthesis (Tsantaki et al.
-2018), to determine the Teff from individual stellar spectra.
-FASMA relies on the radiative transfer code MOOG444https://www.as.utexas.edu/~chris/moog.html (Sneden 1973) to model the atmosphere and create a synthetic spectrum for a set of input parameters (surface gravity, ; iron metallicity, [Fe/H]; macro-turbulence velocity ; and projected rotational velocity ) of a given star.
-The synthetic spectrum is then statistically compared to the observed spectrum of the star and the parameters are iteratively adjusted using a non-linear least-squares algorithm to find the set that best fits observations.
-To capture reliable variations in Teff, when computing it, we fix all other input stellar parameters, and leave Teff to be computed freely for each individual stellar spectra.
+We used the stellar spectral analysis package FASMA555https://github.com/MariaTsantaki/FASMA-synthesis (Tsantaki et al.
+2018) to determine the Teff from individual stellar spectra.
+FASMA relies on the radiative transfer code MOOG666https://www.as.utexas.edu/~chris/moog.html (Sneden 1973) to model the atmosphere and create a synthetic spectrum for a set of input parameters of a given star, which include surface gravity, ; iron metallicity, [Fe/H]; macro-turbulence velocity ; and projected rotational velocity .
+The synthetic spectrum was then statistically compared to the observed spectrum of the star, and the parameters were iteratively adjusted using a non-linear least-squares algorithm in order to find the set that best fits the observations.
+To capture reliable variations in Teff, when computing it, we fixed all other input stellar parameters and left Teff to be computed freely for each individual stellar spectra.
 Stellar parameters for the stars in our sample have been derived in Tsantaki et al.
-(2023) also using FASMA.
+(2023) also by using FASMA.
 4 Results and discussion
 As presented in Table 3 and in accordance with Mermilliod et al.
-(2008), the star NGC3532No649 is not a member of cluster NGC3532 making the calculation of its mass and evolutionary stage uncertain.
-For this reason, it will not be included in the following discussion despite it being Li-rich.
-4.1 Is the Lithium Abundance of Red Giants related to Stellar Activity?
-We start by analysing the dispersion (STD and PTP) of BIS, FWHM, T and H to probe potential connections between activity variability and A(Li).
-We also test if the mean and median of Hα, indicative of a star’s activity level, is related to A(Li).
-For the computation of these relations we only use stars with A(Li) as this is the limit of FASMA’s minimization algorithm in the determination of A(Li) (Tsantaki et al.
+(2008), the star NGC3532No649 is not a member of cluster NGC3532, making the calculation of its mass and evolutionary stage uncertain.
+For this reason, it is not included in the following discussion despite it being Li-rich.
+4.1 The relationship between lithium abundance and stellar activity in red giants
+We started by analysing the dispersion (STD and PTP) of BIS, FWHM, T, and H in order to probe for any potential connections between activity variability and A(Li).
+We also tested if the mean and median of Hα, which are indicative of a star’s activity level, are related to A(Li).
+For the computation of these relations, we only used stars with A(Li), as this is the limit of FASMA’s minimisation algorithm in the determination of A(Li) (Tsantaki et al.
 2023).
 Using these limit values would be a source of bias.
-We find that the STD and PTP of BIS and FWHM increase exponentially with A(Li) and the mean and median values of H06 also increase exponentially with A(Li).
-In particular, the vast majority of stars with high H06 index are Li enriched.
-On the other hand, we do not find A(Li) to be related to H16.
+We find that the STD and PTP of BIS and FWHM increase exponentially with A(Li), and the mean and median values of H06 also increase exponentially with A(Li).
+In particular, the vast majority of stars with a high H06 index are Li enriched.
+On the other hand, we did not find A(Li) to be related to H16.
 These results suggest that the activity levels of the stars in our sample seem to increase with increasing A(Li).
 These results are in agreement with Sneden et al.
-(2022) who found that Li-rich giants are more chromospherically active when in comparison to their Li poor counterparts, as they are more likely to show strong He I10830 absorptions.
-We also find that correlates well with dispersion of BIS and FHWM as well as the mean and median of the H06 index.
+(2022), who found that Li-rich giants are more chromospherically active in comparison to their Li-poor counterparts, as they are more likely to show strong He I10830 absorptions.
+We also find that correlates well with the dispersion of BIS and FHWM as well as the mean and median of the H06 index.
 This is indicative that more magnetically active stars also tend to rotate faster.
@@
 (2020) also arrived at a similar conclusion.
-The best fit relations are listed in Table 3.
-Plots of the relations found for STD of BIS and FWHM as well as the mean of H06 are shown in Figure 1.
-The plots for PTP of BIS and FWHM and the median of H06 look similar to the ones shown in Figure 1 and are presented in Appendix B.
-Also in Appendix B are plots of the mean and median H16 as a function of A(Li), for which no clear relation was found.
-In addition to the results of the F-test, in Table 3 we also provide R2 values, with better fits having R2 closer to 1.
-R2 values for the relations we find are in general low as there is significant scatter around the lines of best fit.
+The best-fit relations are listed in Table 7.
+Plots of the relations found for the STD of BIS and FWHM as well as the mean of H06 are shown in Figure 1.
+The plots for the PTP of BIS and FWHM and the median of H06 look similar to the ones shown in Figure 1 and are presented in Appendix D.
+Also in Appendix D are plots of the mean and median H16 as a function of A(Li), for which no clear relation was found.
+In addition to the results of the F-test, in Table 7 we also provide R2 values, with better fits having R2 closer to one.
+The R2 values for the relations we found are in general low, as there is significant scatter around the lines of the best fit.
 | Formula | F-Statistic | p(F-Statistic) | R2 |
@@
 | = | 34.82 | 1.09 | 0.33 |
-4.2 Manifestation of Stellar Activity in RV signals and its Connection to Lithium Enhancement
+4.2 Manifestation of stellar activity in radial velocity signals and its connection to lithium enhancement
 To further investigate these findings, we also looked at how RV variability correlates with the stellar activity indicators and whether this is linked to A(Li).
-All the correlations discussed here are shown in Figure 2 and listed in Appendix C.
-From our analysis, 22/81 stars show a strong correlation (0.4555Some of the analysed stars in our sample do not have a large number of data points.
-Because of this we chose to include moderate correlations which have the potential to become strong with the addition of new data.
-For ease of nomenclature we refer to all correlations 0.4 as strong.) between the RV variability and the BIS variability.
+All the correlations discussed here are shown in Figure 2 and listed in Appendix B.
+From our analysis, 22/81 stars show a strong correlation (0.4888Some of the analysed stars in our sample do not have a large number of data points.
+Because of this, we chose to include moderate correlations that have the potential to become strong with the addition of new data.
+For ease of nomenclature, we refer to all correlations greater than 0.4 as strong.) between the RV variability and the BIS variability.
 The majority of the correlations were found to be negative.
-We noticed that 13/22 of the strong correlations found between RV and BIS appear for stars that are moderate and fast rotators, that is, show (e.g.
+As stars age, their decreases, typically falling below 5 km/s by the end of the main-sequence (e.g.
 de Medeiros et al.
@@
 2015; Patton et al.
-2023).
+2024).
+We noticed that 13/22 of the strong correlations found between RV and BIS appear for stars that are moderate and fast rotators ().
 This is likely a consequence of this diagnostic of line asymmetry losing sensitivity for slower rotators (e.g.
@@
 2022; Sayeed et al.
-2023; Tsantaki et al.
+2024; Tsantaki et al.
 2023) that there is a connection between Li enrichment in giants and .
-This is also suggested in our sample as we find that increases exponentially with A(Li) (shown in Appendix B and Table 3).
-Out of the 27 Li-rich stars listed in Table 2, 8 stars show strong correlations between RV and BIS.
+This is also suggested in our sample, as we find that increases exponentially with A(Li) (shown in Appendix D and Table 7).
+Out of the 27 Li-rich stars listed in Table 3, eight stars show strong correlations between RV and BIS.
 In what concerns the FWHM, 20/81 stars show strong correlations with RV measurements.
 The correlations are preferentially positive.
-6 of the Li-rich stars in our sample show strong correlations between RV and FWHM.
-For the T we find that 15/81 stars show strong correlations with RV measurements.
+Six of the Li-rich stars in our sample show strong correlations between RV and FWHM.
+For the T, we find that 15/81 stars show strong correlations with RV measurements.
 Correlations between the RV and Teff are mostly positive.
-Only 3 of the Li-rich stars show strong correlations between RV and T.
-There is a possibility that this is such because the individual spectra used to measure Teff values to look for variations may not have a signal to noise ratio high enough for this purpose.
-As for H16, 20/81 stars show strong correlations with RV, which are mostly negative. 9 of the Li-rich stars show strong correlations between RV and H16.
-Last but not least for H06, 19/81 stars show significant correlations with RV measurements, which are preferentially positive.
-We find that 10 of the Li-rich stars in our sample show strong correlations between RV and H06.
+Only three of the Li-rich stars show strong correlations between RV and T.
+There is a possibility that this the case is because the individual spectra used to measure Teff values to look for variations may not have a signal-to-noise ratio high enough for this purpose.
+As for H16, 20/81 stars show strong correlations with RV, and they are mostly negative. Nine of the Li-rich stars show strong correlations between RV and H16.
+Last but not least, for H06, 19/81 stars show significant correlations with RV measurements, and they are preferentially positive.
+We find that ten of the Li-rich stars in our sample show strong correlations between RV and H06.
 Overall, correlations were found for stars in various evolutionary stages and of different mass ranges as well.
-This is evidenced in the Hertzsprung–Russell (HR) diagram Figure 3, where different symbols indicate different numbers of strong correlations found per star, and the data points are sized according to stellar mass.
-In general, all symbols display a range of sizes and are distributed across the HR diagram revealing there is no preferred mass or age for correlations to be found.
-A summary of the significant correlations discussed can be found in Tables 4 and 5.
+This is evidenced in the Hertzsprung–Russell (HR) diagram in Figure 3, where different symbols indicate different numbers of strong correlations found per star and the data points are sized according to stellar mass.
+In general, all symbols display a range of sizes and are distributed across the HR diagram, revealing there is no preferred mass or age for correlations to be found.
+A summary of the significant correlations discussed can be found in Tables 9 and 10.
 Something to note from Table 10 is that inhabitants of cluster NGC6705, in general, show higher amounts of Li than the rest of the stars in our sample.
-These stars are particular as they also demonstrate unusually high amounts of -elements for their young age (e.g.
+These stars are particular, as they also demonstrate unusually high amounts of -elements for their young age (e.g.
 Casamiquela et al.
@@
 Messina et al.
-(2010) found that the G type stars of this cluster retain higher levels of activity up to the age of the cluster.
-We find that, 19/28 of these stars show at least one significant correlation with one of the parameters at study and 22/28 are fast rotators (see Table 5).
-So in general, the stars in this cluster are magnetically active and also rotate fast which could be connected to the abnormally high levels of Li observed.
-In Figure 4 we show a violin plot of A(Li) as a function of the number of strong correlations observed per star.
-We may see that the density of points increases towards higher values of A(Li) with increasing number of correlations found.
-This is evidenced by the increasing width of the violins from left to right in high A(Li) regions and by the central tendency being slightly higher for stars that show 1 or more than 2 correlations between RV and stellar activity indicators.
+(2010) found that the G-type stars of this cluster retain higher levels of activity up to the age of the cluster.
+We find that 19/28 of these stars show at least one significant correlation with one of the parameters we studied, and 22/28 are fast rotators (see Table 10).
+So, in general, the stars in this cluster are magnetically active and also rotate fast, which could be connected to the abnormally high levels of Li observed.
+In Figure 4, we show a violin plot of A(Li) as a function of the number of strong correlations observed per star.
+One may see that the density of the points increases towards higher values of A(Li), with an increasing number of correlations found.
+This is evidenced by the increasing width of the violins from left to right in high A(Li) regions and by the central tendency being slightly higher for stars that show one or more than two correlations between RV and stellar activity indicators.
 As such, strong correlations tend to appear more often for stars with higher amounts of Li in their atmospheres.
 Such a result also indicates that A(Li) seems to be connected with higher levels of stellar activity in our sample.
-To complement the correlation analysis, GLSP of RV and stellar activity indicators were also plotted for all stars.
-GLSP are often used when analysing whether variability observed in RV signals is caused by the existence of a companion or by stellar activity.
-This is because, they allow us to determine if there is periodic variability occurring at the stellar surface that is equivalent to periodic variability observed in RV.
-It has been shown that magnetic phenomena appearing at the stellar surface can introduce variability in RV that is at a different phase relatively to variability observed in stellar activity indicator signals (e.g.
+To complement the correlation analysis, GLSPs of RV and stellar activity indicators were also plotted for all stars.
+The GLSPs are often used when analysing whether variability observed in RV signals is caused by the existence of a companion or by stellar activity.
+This is because they make it possible to determine if there is periodic variability occurring at the stellar surface that is equivalent to the periodic variability observed in RV.
+It has been shown that magnetic phenomena appearing at the stellar surface can introduce variability in RV that is at a different phase relative to the variability observed in stellar activity indicator signals (e.g.
 Queloz et al.
@@
 2018).
-As such, when looking for correlations between RV and activity proxies we may obtain a weak correlation coefficient, but this does not necessarily mean that RV variability does not originate from stellar activity.
-Thus, finding peaks at similar periods in RV and again in different stellar activity signals with GLSP, is also an indicator that RV variability connects with stellar activity.
-The results of the GLSP analysis are also presented in Tables 4 and 5.
-In summary, stars NGC2345No50, NGC2423No3, NGC3532No670, NGC4349No127 and NGC6705No1101 show significant peaks (above a false alarm probability (FAP) level of 1%) in their RV GLSPs which consistently appear around the same period for at least one of their stellar activity indicators.
-All relevant GLSPs are in Appendix D and Delgado Mena et al.
+As such, when looking for correlations between RV and activity proxies, we may obtain a weak correlation coefficient, but this does not necessarily mean that the RV variability does not originate from stellar activity.
+Thus, finding peaks at similar periods in RV and again in different stellar activity signals with GLSP is also an indicator that RV variability connects with stellar activity.
+The results of the GLSP analysis are also presented in Tables 9 and 10.
+In summary, stars NGC2345No50, NGC2423No3, NGC3532No670, NGC4349No127, and NGC6705No1101 show significant peaks (above a false-alarm probability (FAP) level of 1%) in their RV GLSPs, which consistently appear around the same period for at least one of their stellar activity indicators.
+All relevant GLSPs are in Appendix E and Delgado Mena et al.
 (2023).
-All of these stars can be considered Li-rich (see Table 2) and most of them also present long-term and large amplitude RV variations which mimic the presence of planets but are of probable stellar origin (Delgado Mena et al.
+All of these stars can be considered Li-rich (see Table 3), and most of them also present long-term and large amplitude RV variations that mimic the presence of planets but are of probable stellar origin (Delgado Mena et al.
 2023).
@@
 Interestingly, Jorissen et al.
-(2020) studying the binarity fraction of Li-rich giants found that some of the most luminous Li-rich giants show RV variability not caused by stellar companions, though no information about stellar activity is provided by the authors.
-In a similar way, Gonçalves et al.
-(2020) finds hints of RV variability in several Li-rich stars with detected longitudinal magnetic field, but the quantity of data does not allow to discern the real origin of such variability.
+(2020), studying the binarity fraction of Li-rich giants, found that some of the most luminous Li-rich giants show RV variability not caused by stellar companions, though no information about stellar activity is provided by the authors.
+Similarly, Gonçalves et al.
+(2020) found hints of RV variability in several Li-rich stars with a detected longitudinal magnetic field, but the quantity of data does not permit one to discern the real origin of such variability.
 5 Conclusions
-In summary based on the results from HARPS data, we find that the variability of BIS and FWHM and the mean and median of H06 increase exponentially with A(Li).
+In summary, based on the results from the HARPS data, we find that the variability of BIS and FWHM and the mean and median of H06 increase exponentially with A(Li).
 This leads us to conclude that the stars in our sample that are more magnetically active also show higher amounts of Li in their atmosphere.
 We also find that stars with higher amounts of Li tend to show more correlations between RV and stellar activity indicators.
-The majority of stars, 19/27 (10/14 leaving aside the special cluster NGC6705), that can be considered Li-rich in this work, either show correlations of RV with at least one of the stellar activity indicators or show periodic variability in at least one of the stellar activity indicator signals with a similar period in RV.
+The majority of the stars, 19/27 (10/14 leaving aside the special cluster NGC6705), that can be considered Li-rich in this work either show correlations of RV with at least one of the stellar activity indicators or show periodic variability in at least one of the stellar activity indicator signals with a similar period in RV.
 From the GLSP analysis we may also conclude that almost all the stars with strong RV variations that mimic the presence of planets are Li-rich (see also, Delgado Mena et al.
 2016).
-This is particularly important when searching for exoplanets orbiting evolved stars, since our results show that detecting periodic RV variability in red giants with high A(Li) might be a signature of stellar activity instead of planets.
+This is particularly important when searching for exoplanets orbiting evolved stars since our results show that detecting periodic RV variability in red giants with high A(Li) might be a signature of stellar activity instead of planets.
 It is interesting to note that Adamów et al.
-(2018) reported a higher than expected fraction of planetary companions around the Li-rich stars in their sample, pointing to a connection between Li enrichment and planetary presence/engulfment.
-However, the orbits they find are quite eccentric, something not expected if the RV were of stellar origin, and that supports the hypothesis of inner planets engulfment.
-The Li-rich stars with strong RV variability discussed here lie on the RGB or RGB tip, however, stars showing correlations between Li abundance and activity level are found all across the HR diagram.
-This might indicate that mild Li enrichment (since we do not have super-Li-rich giants in our sample) connected to stellar activity can occur at different evolutionary stages, but a large enhancement of Li is more particular and can only occur at certain evolutionary stages, such as, the He-flash (Singh et al.
+(2018) reported a higher than expected fraction of planetary companions around the Li-rich stars in their sample, pointing to a connection between Li enrichment and planetary presence or engulfment.
+However, the orbits they found are quite eccentric, something not expected if the RV is of stellar origin, and that characteristic supports the hypothesis of inner planet engulfment.
+The Li-rich stars with strong RV variability discussed here lie on the RGB or RGB tip.
+However, stars showing correlations between Li abundance and activity level are found all across the HR diagram.
+This might indicate that mild Li enrichment (since we do not have super Li-rich giants in our sample) connected to stellar activity can occur at different evolutionary stages, but a large enhancement of Li is more particular and can only occur at certain evolutionary stages, such as the He flash stage (Singh et al.
 2019; Kumar et al.
 2020).
-It is still not clear then, whether the strong RV variability of mild Li-rich giants on the RGB is connected to their evolutionary stage and the possible super-Li rich phase to happen soon after (at the beginning of the He-core burning phase), or if it is a consequence of stellar activity.
+Thus, it is still not clear whether the strong RV variability of mild Li-rich giants on the RGB is connected to their evolutionary stage and the possible super Li-rich phase that happens soon after (at the beginning of the He-core burning phase) or if it is a consequence of stellar activity.
 Acknowledgements.
+We thank the anonymous referee for the insightful comments and suggestions.
 I.R.
@@
 E.D.M and J.G.S.
-acknowledge the support from FCT through the grant 2022.04416.PTDC.References
+acknowledge the support from FCT through the grant 2022.04416.PTDC.
+This work made use of scientific colour maps (https://www.fabiocrameri.ch/colourmaps/ Crameri et al.
+2020) to mitigate visual distortion.References
 - Adamów et al. (2018) Adamów, M., Niedzielski, A., Kowalik, K., et al. 2018, A&A, 613, A47
@@
 - Aguilera-Gómez et al. (2023) Aguilera-Gómez, C., Jones, M. I., & Chanamé, J. 2023, A&A, 670, A73
-- Bagdonas et al. (2018) Bagdonas, V., Drazdauskas, A., Tautvaišienė, G., Smiljanic, R., & Chorniy, Y. 2018, Astronomy and Astrophysics, 615, A165
+- Bagdonas et al. (2018) Bagdonas, V., Drazdauskas, A., Tautvaišienė, G., Smiljanic, R., & Chorniy, Y. 2018, A&A, 615, A165
 - Brown et al. (1989) Brown, J. A., Sneden, C., Lambert, D. L., & Dutchover, Edward, J. 1989, ApJS, 71, 293
@@
 - Charbonnel (1995) Charbonnel, C. 1995, ApJ, 453, L41
-- Charbonnel & Balachandran (2000) Charbonnel, C. & Balachandran, S. C. 2000, arXiv preprint astro-ph/0005280
+- Charbonnel & Balachandran (2000) Charbonnel, C. & Balachandran, S. C. 2000, A&A, 359, 563
+- Crameri et al. (2020) Crameri, F., Shephard, G. E., & Heron, P. J. 2020, Nature Communications, 11, 5444
 - de Medeiros et al. (1996) de Medeiros, J. R., Melo, C. H. F., & Mayor, M. 1996, A&A, 309, 465
@@
 - Delgado Mena et al. (2018) Delgado Mena, E., Lovis, C., Santos, N. C., et al. 2018, A&A, 619, A2
-- Delgado Mena et al. (2016) Delgado Mena, E., Tsantaki, M., Sousa, S., et al. 2016, Astronomy & Astrophysics, 587, A66
+- Delgado Mena et al. (2016) Delgado Mena, E., Tsantaki, M., Sousa, S. G., et al. 2016, A&A, 587, A66
 - Denissenkov & Herwig (2004) Denissenkov, P. A. & Herwig, F. 2004, ApJ, 612, 1081
@@
 - Fekel & Balachandran (1994) Fekel, F. C. & Balachandran, S. 1994, in Cool Stars, Stellar Systems, and the Sun, Vol. 64, 279
-- Frinchaboy & Majewski (2008) Frinchaboy, P. M. & Majewski, S. R. 2008, The Astronomical Journal, 136, 118
+- Frinchaboy & Majewski (2008) Frinchaboy, P. M. & Majewski, S. R. 2008, AJ, 136, 118
 - Gao et al. (2019) Gao, Q., Shi, J.-R., Yan, H.-L., et al. 2019, ApJS, 245, 33
@@
 - Gonçalves et al. (2020) Gonçalves, B. F. O., da Costa, J. S., de Almeida, L., Castro, M., & do Nascimento, J. D., J. 2020, MNRAS, 498, 2295
-- Herzog et al. (1975) Herzog, A. D., Sanders, W. L., & Seggewiss, W. 1975, Astronomy and Astrophysics Supplement Series, 19, 211
+- Herzog et al. (1975) Herzog, A. D., Sanders, W. L., & Seggewiss, W. 1975, A&AS, 19, 211
 - Holanda et al. (2019) Holanda, N., Pereira, C. B., & Drake, N. A. 2019, MNRAS, 482, 5275
@@
 - Jorissen et al. (2020) Jorissen, A., Van Winckel, H., Siess, L., et al. 2020, A&A, 639, A7
-- Katime Santrich et al. (2013) Katime Santrich, O. J., Pereira, C. B., & Drake, N. A. 2013, Astronomy and Astrophysics, 554, A2
-- Kirby et al. (2016) Kirby, E. N., Guhathakurta, P., Zhang, A. J., et al. 2016, 819, 135
+- Katime Santrich et al. (2013) Katime Santrich, O. J., Pereira, C. B., & Drake, N. A. 2013, A&A, 554, A2
+- Kirby et al. (2016) Kirby, E. N., Guhathakurta, P., Zhang, A. J., et al. 2016, ApJ, 819, 135
 - Kowkabany et al. (2022) Kowkabany, J., Ezzeddine, R., Charbonnel, C., et al. 2022, arXiv e-prints, arXiv:2209.02184
@@
 - Liu et al. (2014) Liu, Y. J., Tan, K. F., Wang, L., et al. 2014, ApJ, 785, 94
-- Lovis & Mayor (2007) Lovis, C. & Mayor, M. 2007, Astronomy and Astrophysics, 472, 657
-- Luck (1994) Luck, R. E. 1994, The Astrophysical Journal Supplement Series, 91, 309
+- Lovis & Mayor (2007) Lovis, C. & Mayor, M. 2007, A&A, 472, 657
+- Luck (1994) Luck, R. E. 1994, ApJS, 91, 309
 - Lyubimkov (2016) Lyubimkov, L. S. 2016, Astrophysics, 59, 411
@@
 - Mayor et al. (2003) Mayor, M., Pepe, F., Queloz, D., et al. 2003, The Messenger, 114, 20
-- Mermilliod et al. (2008) Mermilliod, J. C., Mayor, M., & Udry, S. 2008, Astronomy and Astrophysics, 485, 303
+- Mermilliod et al. (2008) Mermilliod, J. C., Mayor, M., & Udry, S. 2008, A&A, 485, 303
 - Messina et al. (2010) Messina, S., Parihar, P., Koo, J.-R., et al. 2010, A&A, 513, A29
 - Pabst et al. (2020) Pabst, C. H. M., Goicoechea, J. R., Teyssier, D., et al. 2020, A&A, 639, A2
-- Palacios et al. (2004) Palacios, A., Charbonnel, C., Talon, S., & Siess, L. 2004, arXiv e-prints, astro
-- Pallavicini et al. (1992) Pallavicini, R., Randich, S., & Giampapa, M. 1992, Astronomy and Astrophysics, 253, 185
-- Patton et al. (2023) Patton, R. A., Pinsonneault, M. H., Cao, L., et al. 2023, arXiv e-prints, arXiv:2303.08151
+- Palacios et al.
+(2006) Palacios, A., Charbonnel, C., Talon, S., & Siess, L.
+2006, in Chemical Abundances and Mixing in Stars in the Milky Way and its Satellites, ed.
+S.
+Randich & L.
+Pasquini (Berlin, Heidelberg: Springer Berlin Heidelberg), 304–305
+- Pallavicini et al. (1992) Pallavicini, R., Randich, S., & Giampapa, M. S. 1992, A&A, 253, 185
+- Patton et al. (2024) Patton, R. A., Pinsonneault, M. H., Cao, L., et al. 2024, MNRAS, 528, 3232
 - Peña Suárez et al. (2018) Peña Suárez, V. J., Sales Silva, J. V., Katime Santrich, O. J., Drake, N. A., & Pereira, C. B. 2018, ApJ, 854, 184
 - Queloz et al. (2001) Queloz, D., Henry, G. W., Sivan, J. P., et al. 2001, A&A, 379, 279
-- Queloz et al. (2000) Queloz, D., Mayor, M., Weber, L., et al. 2000, Astronomy and Astrophysics, 354, 99
+- Queloz et al. (2000) Queloz, D., Mayor, M., Weber, L., et al. 2000, A&A, 354, 99
 - Randich et al. (2020) Randich, S., Pasquini, L., Franciosini, E., et al. 2020, A&A, 640, L1
@@
 - Santos et al. (2003) Santos, N. C., Udry, S., Mayor, M., et al. 2003, A&A, 406, 373
-- Sayeed et al. (2023) Sayeed, M., Ness, M. K., Montet, B. T., et al. 2023, arXiv e-prints, arXiv:2306.03323
+- Sayeed et al. (2024) Sayeed, M., Ness, M. K., Montet, B. T., et al. 2024, ApJ, 964, 42
 - Singh et al. (2019) Singh, R., Reddy, B. E., Bharat Kumar, Y., & Antia, H. M. 2019, ApJ, 878, L21
@@
 - Sneden et al. (2022) Sneden, C., Afşar, M., Bozkurt, Z., et al. 2022, ApJ, 940, 12
-- Sneden (1973) Sneden, C. A. 1973, PhD thesis, University of Texas
-- Sun et al. (2022) Sun, Q., Deliyannis, C. P., Twarog, B. A., et al. 2022, Monthly Notices of the Royal Astronomical Society, 513, 5387
+- Sneden (1973) Sneden, C. A. 1973, PhD thesis, The University of Texas at Austin, USA
+- Sun et al. (2022) Sun, Q., Deliyannis, C. P., Twarog, B. A., et al. 2022, MNRAS, 513, 5387
 - Tayar et al. (2015) Tayar, J., Ceillier, T., García-Hernández, D. A., et al. 2015, ApJ, 807, 82
@@
 - Tsantaki et al. (2023) Tsantaki, M., Delgado-Mena, E., Bossini, D., et al. 2023, A&A, 674, A157
-- Udry et al. (2000) Udry, S., Mayor, M., Naef, D., et al. 2000, Astronomy and Astrophysics, 356, 590
+- Udry et al. (2000) Udry, S., Mayor, M., Naef, D., et al. 2000, A&A, 356, 590
 - Xing et al. (2021) Xing, L.-F., Li, Y.-C., Chang, L., Wang, C.-J., & Bai, J.-M. 2021, A&A, 653, A28
 - Yana Galarza et al. (2019) Yana Galarza, J., Meléndez, J., Lorenzo-Oliveira, D., et al. 2019, MNRAS, 490, L86
-- Zboril et al. (1997) Zboril, M., Byrne, P. B., & Rolleston, W. R. J. R. 1997, Monthly Notices of the Royal Astronomical Society, 284, 685
+- Zboril et al. (1997) Zboril, M., Byrne, P. B., & Rolleston, W. R. J. R. 1997, MNRAS, 284, 685
 - Zechmeister & Kürster (2009) Zechmeister, M. & Kürster, M. 2009, A&A, 496, 577
+Appendix A Additional tables
 | Star | (RV,BIS) | (RV,FWHM) | (RV,Teff) | (RV,H1.6Å) | (RV,H0.6Å) | A(Li) [dex] | Mass (M⊙) | (km/s) | n |
@@
 | NGC6705No963 | 1.542 | 3.757 | 6.710 | 9 |
-Appendix A Methods for Data Analysis
-Appendix B Statistically Significant Relations
-Appendix C Correlation Tables
+Appendix B Correlation tables
 | Star | (RV,BIS) | p(BIS) | (RV,FWHM) | p(FWHM) | (RV,Teff) | p(Teff) | (RV,H1.6Å) | p(H1.6Å) | (RV,H0.6Å) | p(H0.6Å) | A(Li) [dex] | Mass (M⊙) | (km/s) | n |
@@
 | NGC6705No963 | -0.28 | 0.47 | 0.27 | 0.48 | 0.03 | 0.93 | -0.21 | 0.60 | 0.14 | 0.71 | 1.542 | 3.757 | 6.710 | 9 |
-Appendix D Generalized Lomb-Scargle Periodograms
+Appendix C Methods for data analysis
+Appendix D Additional relations
+Appendix E Generalized Lomb-Scargle periodograms
 The remainder of the GLPs mentioned in the text can be found in Delgado Mena et al. (2023).